### v1

# Optimizing Multi-Modal Trackers via Sensitivity-aware Regularized Tuning

Abstract
This paper tackles the critical challenge of optimizing multi-modal trackers by effectively adapting the pre-trained models for RGB data. Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off. To mitigate this dilemma, we propose a novel sensitivity-aware regularized tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities. Through a comprehensive investigation from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are primary drivers of this issue. Specifically, we first analyze the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization. Then, we further explore transfer sensitivities during the tuning phase, emphasizing adaptability and stability. By incorporating these sensitivities as regularization terms, our method significantly enhances the transferability across modalities. Extensive experiments showcase the superior performance of the proposed method, surpassing current state-of-the-art techniques across various multi-modal tracking. The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
Index Terms:
Multi-modal Tracking, Cross-modal Transfer, Parameter Sensitivity, Regularized Tuning.
Figure 1: Training and testing performance on LasHeR, comparing our method with existing fine-tuning approaches. Our method effectively mitigates the misfitting issue and enhances both generalization and stability of the multi-modal tracker.
I Introduction
Object tracking, a foundation task of visual perception, has seen significant advancements over the past decades [1, 2, 3, 4]. Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions. Therefore, multi-modal tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest. With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [5, 6, 7, 8]. There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
Some existing approaches have explored the full fine-tuning (FFT) paradigm [9, 10, 11], where these trackers are initialized with pre-trained weights and subsequently tuned for task-specific objectives. Nevertheless, due to the significant gap and limited scale of auxiliary modalities, they are intractable to retain the pre-trained knowledge during transfer and tend to induce catastrophic forgetting and overfitting (i.e., widened train-test gap). In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [12, 13, 14]. The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modal-specific ones to retain pre-trained knowledge. Several methods fall under this umbrella [15, 16, 17], including prompt tuning, visual adapter, etc. While effective, PEFT methods impose excessive constraints on the primary model weights, resulting in underfitting (i.e., restricted train upper bound) for handling the vast distribution drift. In summary, existing tuning paradigms fluctuate between excessive flexibility and overly rigid constraints, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and adaptation (please refer to Fig. 1).
In this work, we endeavor to mitigate the misfitting issue in cross-modal tracker adaptation by strategically regularizing the learning process. By examining parameter responses spanning from the pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability. To mitigate this, we propose a sensitivity-aware regularized tuning framework that delicately calibrates the gradient updates to boost model transfer with precision. Specifically, we optimize multi-modal trackers from the following perspectives.
  1. (1)
    Formulating Prior Sensitivity. We begin by investigating the tangent space of pre-trained parameters, using it as a critical indicator to assess and preserve prior generalizability. Moreover, we estimate such prior sensitivity via an eigen-decomposition approximation.
  2. (2)
    Modeling Transfer Sensitivity. To further elucidate transfer challenges, we explore how sparse gradients exacerbate tuning instability. Accordingly, we construct this transfer sensitivity using the off-the-shelf gradient matrix, with the aim of facilitating gradient rebalancing.
  3. (3)
    Sensitivity-Aware Regularization Tuning. By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modal representations through finely modulated, sensitivity-driven updates. This regularization tuning facilitates seamless transfer to various multi-modal tracking tasks, which can continuously enhance the model in the training phase.
Our method strategically guides the cross-modal fine-tuning process to optimize downstream tracking tasks. Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modal tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models. Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
In summary, the main contributions of this paper are:
  • we revisit the misfitting issue of multi-modal tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of multi-modal trackers;
  • we conduct comprehensive experiments covering three multi-modal tracking tasks and seven datasets using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels;
The remainder of the paper is organized as follows. Section II reviews existing works on multi-modal trackers and cross-modal transfer learning methods. In Section III, we explicitly formulate the prior and transfer sensitivity from both the pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-aware regularized tuning. In Section IV, we present experiments on seven multi-modal tracking benchmarks to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs. Finally, Section V concludes the paper.
II Related Work
II-A Multi-Modal Tracking
Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [18, 19, 20, 21]. In recent years, due to the unreliability of RGB data in challenging scenarios, increasing studies are expected to integrate auxiliary modalities (e.g., event, depth, thermal data) to enhance tracking performance. These complementary signals compensate for the inherent limitations of RGB imagery [22]. For instance, event cameras provide robust dynamic scene information under extreme motion or lighting. Leveraging this, [23, 24, 25, 26, 27, 28] combine RGB and event streams to predict objects in high-speed and low-dynamic scenarios. Meanwhile, depth data offers spatial cues that are particularly useful for handling occlusion and background clutter. [29, 30, 31, 32] exploit this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes. Similarly, thermal infrared imaging captures heat signatures independent of visible light, proves effective in low-illumination. [33, 34, 35, 36] demonstrate that fusing thermal and RGB data can yield more reliable appearance representations.
In summary, these methods emphasize effective multi-modal interaction and fusion [37, 9, 38, 39]. With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [5, 40, 8] have achieved remarkable generalization across diverse scenarios. These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers. Multi-modal tracking is increasingly driven by reusing RGB pre-trained semantic priors, particularly when annotated multi-modal data are scarce or noisy. In this work, we target this challenge by focusing on optimizing the adaptation of pre-trained trackers for efficient cross-modal transfer.
II-B Cross-modal Transfer Learning
To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged. Some works follow the full fine-tuning (FFT) paradigm [37, 9, 10, 11], where the pre-trained models serve as initializations and are entirely re-trained on the target tasks. These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model. Representatively, SUTrack [8] employs a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization. While often effective, one primary drawback may be innate to this paradigm: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation often leads to significant catastrophic forgetting or overfitting. To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [13, 14, 12], some parameter-efficient fine-tuning (PEFT) methods have been proposed. The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task. By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting. For instance, ProTrack [41] and ViPT [15] modulate RGB features by introducing trainable auxiliary tokens into attention layers. Similarly, methods like BAT [42] and SDSTrack [16] insert lightweight adapter modules between attention layers for cross-modal shift compensation. More recently, unified trackers like UnTrack [17] and OneTracker [1] employ LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs. Despite their effectiveness, excessive constraints on pre-trained weights can incur insufficient adaptation. Furthermore, the added modules tend to over-fit since the additional parameters are optimized from scratch when tailored to modality-specific tasks. Consequently, how to efficiently fine-tune RGB-based pre-trained models for multi-modal tracking remains a significant challenge.
Figure 2: Model architecture for our multi-modality tracker.
III Proposed Method
Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers. To unlock the full potential of pre-trained trackers, we revisit the core principles of cross-modal transfer process. In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning. Next, in Section III-B, we identify pronounced prior and transfer sensitivities that contribute to catastrophic forgetting and training instability, and further quantify these critical sensitivities. Finally, in Section III-C, we incorporate these sensitivities into the tuning phase, enabling adaptive and dynamic modulation of parameter updates.
III-A Preliminaries
Network Architecture. Fig. 2 depicts the overall architecture of our multi-modal tracker. The RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens. Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction. Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens. To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
Cross-Domain Transfer. We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameter $\theta$. It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box. In this paper, we focus on efficiently transferring the models pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$. In this setting, let $\theta_{0}$ be the pre-trained parameters (optimal on $D_{0}$), which serve as initialization. After fine-tuning on $D_{t}$, the parameters become $\theta_{t}$. Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
$\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\theta^{(0)}=\theta_{0},$ (1)
However, directly optimizing $\theta_{t}$ as per Eq. (1) leads to severe catastrophic forgetting and unstable tuning, as evidenced by the pronounced over-fitting and oscillations in training dynamics observed in Fig. 1. This motivates a more principled transfer learning approach described next.
Figure 3: Loss-parameter manifold schematic for different fine-tuning paradigms. The “FFT” fine-tunes all pre-trained weights without regularization, leading to significant catastrophic forgetting on pre-trained dataset. Although “PEFT”-based methods only tune the additional parameters to maintain the pre-traiend knowledge, but their coarse modeling limits performance on new domain. In contrast, our “SRFT” method carefully optimize the model within an approximated tangent space of the pre-trained manifold, indicating a better plasticity-stability trade-off between pre-trained knowledge and adaptation.
III-B Modeling Sensitivity for Multi-Modal Tracking
Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) [43]. We regulate this trade-off via two intrinsic parameter sensitivities, identifying the evolving parameter importance from pre-trained to target tasks: prior sensitivity, reflecting its importance to pre-trained knowledge, and transfer sensitivity, indicating its role in target-task adaptation. Excessive changes to highly sensitive parameters in either aspect can hinder the transfer process.
Manifold hypothesis. As shown in Fig. 3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$. Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$.
Based on the aforementioned manifold hypothesis, to avoid the losing of generalization ability for fine-tuned models on downstream domains, the critical thing lies on optimize the model on local tangent space, shown as the red arrow of the Fig. 3. Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
III-B1 Safeguarding Generalization via Prior Sensitivity
In this context, we first analyze the influence of the tangent space on model generalization from the pre-trained loss-parameter manifold, and then design an efficient eigen-decomposition method to quantify this sensitivity.
Exploring Parameter Manifold Geometry. We first define the joint empirical risks for tuning models over the union of pre-training and downstream tasks as
$\mathcal{L}\left(\theta\mid D_{u}\right)=\mathcal{L}\left(\theta\mid D_{t}\right)+\beta\mathcal{L}\left(\theta\mid D_{0}\right),$ (2)
where $D_{u}=D_{0}\bigcup D_{t}$, $\beta>0$ balances the contributions of pre-trained task, and $\mathcal{L}\left(\theta\mid D_{0}\right)$ serves as a regularization term to refine the vanilla objective. Under this joint objective, the pre-trained weights $\theta_{0}$ minimize $\mathcal{L}\left(\theta\mid D_{0}\right)$ and can serve as a local optimum of $\mathcal{L}\left(\theta\mid D_{u}\right)$. When $\theta$ is in the vicinity of $\theta_{0}$, the pre-training loss can be locally approximated by a second-order Taylor expansion:
$\mathcal{L}\left(\theta\mid D_{0}\right)$ $=\mathcal{L}\left(\theta_{0}\mid D_{0}\right)+\left(\theta-\theta_{0}\right)^{T}\nabla\mathcal{L}\left(\theta_{0}\right)+$ (3)
$\frac{1}{2}\left(\theta-\theta_{0}\right)^{T}\mathcal{F}^{\left(\theta_{0}\right)}\left(\theta-\theta_{0}\right)+\mathcal{O}\left(\left\|\theta-\theta_{0}\right\|^{3}\right),$
where $\nabla\mathcal{L}\left(\theta_{0}\right)\approx\mathbf{0}$, and $\mathcal{F}^{\left(\theta_{0}\right)}\in\mathbb{R}^{|\theta|\times|\theta|}$ is expectation of Hessian matrix $\frac{\partial^{2}\mathcal{L}}{\partial\theta_{i}\theta_{j}}$ over dataset $D_{0}$, also called as the Fisher Information Matrix (FIM) [44] at $\theta_{0}$. Principally, it captures the pre-trained tangent space, along which the impact on the pre-trained loss is minimized.
Eq. (3) reveals that the increase in pre-training loss introduced by transferring weights can be interpreted as the generalization gap:
$\varepsilon_{gen}$ $=\mathcal{L}\left(\theta\mid D_{0}\right)-\mathcal{L}\left(\theta_{0}\mid D_{0}\right)$ (4)
$\approx\frac{1}{2}\left(\theta-\theta_{0}\right)^{T}\mathcal{F}^{\left(\theta_{0}\right)}\left(\theta-\theta_{0}\right)=\frac{1}{2}\left\|\theta-\theta_{0}\right\|_{\mathcal{F}^{\left(\theta_{0}\right)}}^{2},$
This generalization gap is measured as the weight distance within the Riemannian manifold defined by the FIM. Through geometric perspective, the FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [45, 46]. Therefore, the FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
Figure 4: The operation-wise eigen-decomposition FIM map (i.e., prior sensitivity) of the pre-trained OSTrack. The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
Eigen-decomposition FIM Approximation. Directly computing the FIM for a large model is computationally intractable due to its $\mathcal{O}\left(|\theta|^{2}\right)$ complexity. Instead, we propose constructing an eigen-decomposition approximation of the FIM without explicitly forming it in full. Focusing on leading eigenvalues captures key tangent geometrics of affecting generalization while reducing complexity. Specifically, we partition the model’s parameters into $N$ disjoint operation groups, $\theta=\left\{\theta^{1},\cdots,\theta^{N}\right\}$ (e.g. MLPs of FFNs and attention QKVs). Assuming inter-operation independence, we yield a group-diagonal FIM: $\mathcal{F}^{(\theta_{0})}=\mathrm{diag}(\mathcal{F}^{(\theta_{0}^{1})},\dots,\mathcal{F}^{(\theta_{0}^{N})})$, where $\mathcal{F}^{(\theta_{0}^{j})}\in\mathbb{R}^{|\theta^{j}|\times|\theta^{j}|}$ corresponds an independent parameter group $j$. Parameters from different groups thus span orthogonal subspaces with no second-order coupling.
Empirical studies [47, 48] have shown that the FIM of deep networks are spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs. Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
We obtain these leading eigenpairs via Rayleigh-quotient probing [49]. For any non-zero direction $\epsilon$, the Rayleigh quotient is defined as:
$\mathcal{R}(\mathcal{F}^{\left(\theta_{0}\right)},\epsilon)=\frac{\epsilon^{T}\mathcal{F}^{\left(\theta_{0}\right)}\epsilon}{\epsilon^{T}\epsilon},$ (5)
By Rayleigh’s theorem, $\mathcal{R}(\mathcal{F}^{\left(\theta_{0}\right)},\epsilon)$ is maximized when $\epsilon$ aligns with the eigenvector corresponding to the largest eigenvalue of $\mathcal{F}^{\left(\theta_{0}\right)}$. Hence, searching for the “most misdirected” tangents is equivalent to finding the dominant eigenpair $(\lambda_{\max},v_{\max})$, in which case $\lambda_{\max}=\frac{v_{max}^{T}\mathcal{F}^{\left(\theta_{0}\right)}v_{max}}{v_{max}^{T}v_{max}}$. This formula links the sharpest directions of generalization degradation to the dominant FIM eigenvalues, providing both theoretical motivation and practical guidance.
Proposition 1 (Eigen-based Approximation Error Bound)
Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition of group $j$, with top-$K$ eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ and eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$. We construct the following approximation:
$\tilde{\mathcal{F}}^{(\theta_{0}^{j})}=\gamma^{j}I_{|\theta^{j}|},\quad\text{with}\quad\gamma^{j}=\frac{1}{K}\sum_{i=1}^{K}\lambda_{i}^{j},$ (6)
where $\tilde{\mathcal{F}^{(\theta_{0})}}=\mathrm{diag}(\tilde{\mathcal{F}}^{(\theta_{0}^{1})},\dots,\tilde{\mathcal{F}}^{(\theta_{0}^{N})})$ denotes the approximate group-diagonal low-rank FIM. Then the following guarantees hold:
  1. 1.
    Bounded FIM Error. The group-level low-rank approximation $\tilde{\mathcal{F}}^{(\theta_{0})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0})}$ with bounded error. In particular, the Frobenius-norm error satisfies: $\|\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})}\|_{F}\leq\sqrt{\sum_{j=1}^{N}\sum_{i=K+1}^{|\theta^{j}|}{(\lambda_{i}^{j})}^{2}}$, representing the lost Fisher information. If the top-$K$ eigenvalues dominate, $\tilde{\mathcal{F}}^{(\theta_{0})}$ is a close approximation to $\mathcal{F}^{(\theta_{0})}$.
  2. 2.
    Bounded Generalization Gap Error. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded: $|\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})-\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})|\leq\frac{1}{2}\|\Delta\theta\|^{2}(\underset{1\leq j\leq N}{\operatorname{\max}}\lambda_{1}^{j})$.
Proof 1
See Appendix.
Prior Sensitivity Measure. Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [50]. Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to Rayleigh quotient of FIM. This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters. For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
$\lambda^{j}$ $=\underset{\|\epsilon^{j}\|_{2}=\rho\left\|\theta_{0}^{j}\right\|_{2}}{\operatorname{\arg\max}}\frac{(\epsilon^{j})^{T}\mathcal{F}^{\left(\theta_{0}^{j}\right)}(\epsilon^{j})}{(\epsilon^{j})^{T}(\epsilon^{j})}$ (7)
$\approx\underset{\|\epsilon^{j}\|_{2}=\rho\left\|\theta_{0}^{j}\right\|_{2}}{\operatorname{\arg\max}}\frac{\mathcal{L}(\theta_{0}^{j}+\epsilon^{j})-2\mathcal{L}(\theta_{0}^{j})+\mathcal{L}(\theta_{0}^{j}-\epsilon^{j})}{\|\epsilon^{j}\|_{2}^{2}},$
where $\mathcal{L}(\theta_{0}^{j})\triangleq\mathcal{L}(\theta_{0}^{j}|D_{0})$, $\rho=1\times 10^{-5}$ is a noise radius. By repeating this search, we obtain an approximation of the top-$K$ eigenvalues. This measure provides a principled yet tractable perspective to assess how each group of the model affects generalization. By considering the eigenvalue magnitudes from different perturbation directions, we condense the prior sensitivity of the $j^{\text{th}}$ operation into a single scalar by distilling its dominant values:
$s_{j}^{p}=\frac{1}{K}\sum_{i=1}^{K}\lambda_{i}^{j}.$ (8)
The prior sensitivity is equivalent to FIM in Eq. (6), as presented in Fig. 4. We set $K=10$ in our experiments. In practice, we observe that the leading eigenvalue alone (e.g., $K=1$) provides a reliable indicator of prior sensitivity. Further qualitative results are reported in Section IV-C.
III-B2 Stabilizing Adaptation via Transfer Sensitivity
In this context, we analyze the impact of sparse transfer gradients on model adaptation and propose rebalancing them using the gradient matrix.
Characterizing Transfer-Gradient Sparsity. While prior sensitivity addresses forgetting of pre-trained domain, inadaptability in the downstream domain is notably pronounced. Multi-modal tracking often involves diverse and heterogeneous samples exhibiting substantial gaps, such as auxiliary inputs or degraded images. These shifts manifest as highly sparse gradients during fine-tuning process on $D_{t}$, as illustrated in Fig. 5. High sparsity indicates that only a few gradients dominate the updates, while the majority remain nearly static. This phenomenon sheds light on a key factor underlying the limited adaptability of multi-modal trackers and potentially results in temporal oscillatory (see FFT test branch in Fig. 1). In this context, we investigate the elevated adaptation risk induced by sparse gradients and propose a transfer sensitivity aimed at mitigating its impact.
Formally, we quantify the sparsity of the gradient by comparing its $L_{1}$ and $L_{2}$ norms. Let $\mathcal{G}\doteq\nabla_{\theta}\mathcal{L}\left(\theta\mid M\right)$ be the gradient on the current batch $M\subset D_{t}$. We define the sparsity measure: $\rho=\frac{\left\|\mathcal{G}\right\|_{1}}{\sqrt{|\theta|}\left\|\mathcal{G}\right\|_{2}}$. This definition leverages the fact that $L_{1}$ and $L_{2}$ norms characterize different aspects of the vector: while $L_{1}$ reflects the total variation, $L_{2}$ emphasizes concentration [51]. Thus, their ratio serves as an effective proxy for sparsity structure, where a smaller value means higher sparsity. Moreover, noting that single-step gradient magnitudes are typically bounded [52], increasing sparsity (i.e., $\rho:1\rightarrow\frac{1}{\sqrt{|\theta|}}$) can amplify the $L_{2}$ norm:
$\left\|\mathcal{G}\right\|_{2}=\frac{\left\|\mathcal{G}\right\|_{1}}{\rho\sqrt{|\theta|}}.$ (9)
Figure 5: The instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning. The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
Sparsity-Induced Adaptation Risk. We characterize the tuning volatility as the degree to which performance responds to small parameter perturbations during fine-tuning. Accordingly, we define the following sensitivity function:
$S\left(\theta,\delta\mid M\right)=\mathcal{L}\left(\theta\mid M\right)-\mathcal{L}\left(\theta+\delta\mid M\right),$ (10)
where $\delta$ is a random perturbation centered on the current update step, i.e., $\delta\sim\mathcal{N}\left(\alpha\mathcal{G},\sigma^{2}I\right)$, $\alpha$ is the learning rate and $\sigma$ controls the noise. This denotes that updates follow the gradient with inherent noise (due to stochastic sampling, batch gap, etc.).
Intuitively, a smooth and stable loss landscape should ensure that varying perturbations yield consistent responses. The expected magnitude of this difference defines the adaptation risk:
$\varepsilon_{ada}=\underset{\left(\delta,\delta^{\prime}\right)}{\mathbb{E}}\left[|S\left(\theta,\delta\mid M\right)-S\left(\theta,\delta^{\prime}\mid M\right)\right|].$ (11)
By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq. (11), it can be shown that this risk is proportional to the $L_{2}$ norm of gradient:
$\varepsilon_{ada}\approx\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk. This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
Transfer Sensitivity Estimation. To formalize the estimation of transfer sensitivity, we adapt the formulation of Eq. (10) to the standard tuning process. In this case, we adjust the parameter $\theta$ via the gradient $\mathcal{G}$ rather than noise perturbations. Therefore, we can derive: $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\mathcal{G}$. Here, the sensitivity function is used to assess the overall adjustment of all parameters during transfer. To further investigate and regularize gradients, we extend the sensitivity to a parameter-wise formulation:
$s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}})^{2},$ (13)
where $s_{n}^{t}$ denotes the transfer sensitivity of the $n^{\text{th}}$ parameter. This formulation allows us to examine the granular impact of parameter changes during the transfer process and rebalance the gradient accordingly.
1Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
2Output: Optimal multi-modal tracker parameters $\theta_{t}$;
3Step1: Prior Sensitivity Estimation:
for batch $(x,y)$ in $D_{0}$ do
  for $j\in\{1,\ldots,N\}$ do
   Sample $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$;
   Calculate $\lambda^{j}$ via Eq. (7);
   Update eigenvalue set $\Lambda$ for top-$K$ $\lambda^{j}$;
  end for
end for
4 Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
5Step2: Transfer Sensitivity Estimation and Regularized Tuning:
for $i\in\{1,\ldots,T\}$ do
  Sample $i$-th batch data $M_{i}$ from $D_{t}$;
6  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
7  Update transfr sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
8  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
9  Parameter update via Eq. (15);
end for
Algorithm 1 Sensitivity-Aware Regularization Tuning
III-C Sensitivity-Aware Regularization Tuning
We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers. To align with the unified regularization tuning, we rank both the operation-wise prior sensitivity and parameter-wise transfer sensitivity, and then normalize them within the continuous range of $[0,1]$. To harmonize these two sensitivities, we devise a dynamic linear schedule to adjust their weighted combination. At the beginning of training, prior sensitivity plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$). As training progresses, the influence of transfer sensitivity gradually increases, reaching the same weight $\kappa$ by the end. This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability. Formally, at each training step $t$, the combined sensitivity is updated as follows:
$s_{n}=(\kappa+(1-2\kappa)\frac{t}{T})s_{j}^{p}+(1-\kappa-(1-2\kappa)\frac{t}{T})s_{n}^{t},$ (14)
where $\theta_{n}\in\theta^{j}$, $T$ is the total number of training steps. Finally, we normalize this joint sensitivity within the range of $s_{n}\in[0,0.99]$. To this end, during the training process, parameters that are excessively updated will be penalized based on their sensitivity:
$\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\text{s.t. }\theta^{(0)}=\theta_{0},$ (15)
where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments. The detailed regularization tuning process is given in Algorithm.1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Tab. VI.
III-D Learning Objectives
The overall loss function of our method is the same as the foundation model without extra adjustments, shown as:
$\mathcal{L}=\mathcal{L}_{\text{cls}}+\lambda_{\text{iou }}\mathcal{L}_{\text{iou }}+\lambda_{l_{1}}\mathcal{L}_{1},$ (16)
where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as [5].
Figure 6: Visualization of the precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
Figure 7: Precision scores of different attributes on the VisEvent, CoeSot and LasHeR test sets. Zoom in to see details.
IV Experiments
IV-A Experimental Settings
IV-A1 Benchmark Datasets
To verify the effectiveness and generalization of the proposed method, we conduct comprehensive experiments on seven multi-modal benchmark datasets. For RGB-Event tracking, we use FE108 [53] (76 train / 32 test sequences) captured under degraded conditions, VisEvent [9](500 train / 320 test sequences) covering dynamic outdoor scenes, and CoeSot [37] with 578K image-event pairs (824 train / 528 test sequences). For RGB-Depth tracking, we adopt DepthTrack [31] (152 train / 50 test videos) and VOT-RGBD2022 [54] (127 test sequences). For RGB-Thermal tracking, we use LasHeR [55] (979 train / 245 test sequences) and RGBT234 [56] (234 test videos). All training and evaluation strictly follow established protocols [15, 17]. During fine-tuning, we use the training sets of FE108, VisEvent, and CoeSot for RGB-Event tasks; DepthTrack for RGB-Depth; and LasHeR for RGB-Thermal. The corresponding test sets, along with VOT-RGBD2022 and RGBT234, are used for evaluation.
IV-A2 Evaluation Metrics
Adhering to recognized standards [16, 17, 1, 11], we evaluated tracking performance with following metrics. For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR). For DepthTrack, precision (Pr) and recall (Re) are used, with F-score ($F=\frac{2Re\cdot Pr}{Re+Pr}$) as the primary measure. For VOT-RGBD2022, we assess accuracy (Acc), robustness (Rob), and expected average overlap (EAO). For LasHeR, success rate (SR) and precision rate (PR) are employed. For RGBT234, maximum success rate (MSR) and maximum precision rate (MPR) served as evaluation metrics. Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack. For all metrics, the larger, the better.
IV-A3 Pre-trained Models and Compared Methods
In this paper, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack [5], the most widely adopted model; DropTrack [40] and SUTrack [8], which offer improved generalization. Notably, OSTrack and DropTrack are built upon ViT-B/16 [57], while SUTrack adopts the HiViT-B/16 [58] as the backbone. Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack and SUTrack ($\times192$$\times384$) for template and search regions, respectively. To comprehensively validate the effectiveness of our method, we conduct the following experiments in Tab. I and Tab. II. First, we construct strong RGB-based single-modal baselines under full fine-tuning. Then, we evaluate our method under both full and parameter-efficient fine-tuning paradigms. For fair comparison, all methods are grouped according to the pre-trained trackers. Notably, OneTracker [1] and MamTrack [11] adopt non-standard experimental settings and lack open-source implementations. OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Tab. III. MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
IV-A4 Training Details
We follow the data processing pipeline of SDSTrack [16] across all datasets, converting event data into color-polar event images, with no preprocessing for other auxiliary modalities. The models are trained on 8 NVIDIA 3090Ti GPUs with a batch size of 192 and 30 epochs. Each epoch involves sampling 80k samples. We utilize the AdamW optimizer with a learning rate set to $1\times 10^{-4}$ and a weight decay set to $10^{-4}$.
TABLE I: Quantitative comparison on the RGB-Event datasets. The best are marked with “bold”, and the second best are marked with “underline”.
Method Reference Base Model FE108 VisEvent CoeSot
SR PR SR PR SR PR
Image-based Methods (Only-RGB)
DiMP [59] ICCV’19 ResNet 48.5 71.7 45.2 66.2 56.7 69.1
Stark-S [60] ICCV’21 ResNet 51.4 75.1 44.6 61.3 55.7 66.7
OSTrack [5] ECCV’22 OSTrack-256 50.7 73.7 53.4 69.6 64.5 77.0
DropTrack [40] CVPR’23 DropTrack-384 52.0 75.9 54.8 72.4 66.1 79.1
Cross-modal Transfer Learning
FENet [53] ICCV’21 DiMP 63.3 92.4 50.8 70.4 53.8 66.1
AFNet [61] CVPR’23 DiMP 63.8 92.6 51.8 70.8 52.5 63.8
CEUTrack [37] Arxiv’22 OSTrack-B256 56.7 84.2 55.0 71.6 62.0 76.4
ProTrack [41] ACM MM’22 OSTrack-B256 58.0 85.5 47.1 63.2 66.1 78.4
HRTrack [10] ICCV’23 OSTrack-B256 - - - - 63.2 71.9
ViPT [15] CVPR’23 OSTrack-B256 65.1 93.5 59.2 75.8 67.7 80.6
SDSTrack [16] CVPR’24 OSTrack-B256 60.0 87.2 59.7 76.7 67.9 80.9
UnTrack [17] CVPR’24 OSTrack-B256 - - 58.9 75.6 - -
Ours - OSTrack-B256 67.5 96.5 61.8 78.6 69.6 83.0
ViPT [15] CVPR’23 DropTrack-B384 65.6 93.9 60.0 77.4 68.6 81.4
SDSTrack [16] CVPR’24 DropTrack-B384 65.6 94.4 60.6 78.0 69.0 82.2
UnTrack [17] CVPR’24 DropTrack-B384 - - 60.0 77.2 - -
Ours - DropTrack-B384 68.7 97.1 63.0 80.3 71.3 84.9
MamTrack [11] CVPR’25 HiViT-B384 66.4 94.2 61.6 79.2 - -
SUTrack [8] AAAI’25 SUTrack-B384 - - 63.1 79.6 - -
Ours - SUTrack-B384 70.8 98.2 65.4 81.5 73.5 85.6
TABLE II: Quantitative comparison on the RGB-Depth and RGB-Thermal datasets. The best are marked with “bold”, and the second best are marked with “underline”.
Method Reference Base Model DepthTrack VOT RGBD2022 LasHeR RGBT234
Pr Re F-score Acc Rob EAO SR PR MSR MPR
Image-based Methods (Only-RGB)
DiMP [59] ICCV’19 ResNet 46.3 42.8 44.5 70.3 73.1 54.3 42.8 54.4 42.1 62.5
Stark-S [60] ICCV’21 ResNet 39.3 37.6 38.4 65.4 62.8 48.2 37.7 46.3 48.9 66.5
OSTrack [5] ECCV’22 OSTrack-B256 53.6 52.2 52.9 80.3 83.3 67.6 47.2 58.5 54.9 72.9
DropTrack [40] CVPR’23 DropTrack-B384 56.4 55.8 56.1 81.5 85.1 69.2 47.7 59.4 57.2 75.8
Cross-modal Transfer Learning
SPT [62] AAAI’23 Stark-S 52.7 54.9 53.8 79.8 85.1 65.1 33.4 46.7 55.5 78.6
APFNet [63] AAAI’22 Stark-S 51.6 51.4 51.5 79.0 83.7 64.2 36.2 50.0 57.9 82.7
ProTrack [41] ACM MM’22 OSTrack-B256 58.3 57.3 57.8 80.1 80.2 65.1 42.0 53.8 59.9 79.5
ViPT [15] CVPR’23 OSTrack-B256 59.2 59.6 59.4 81.5 87.1 72.1 52.5 65.1 61.7 83.5
SDSTrack [16] CVPR’24 OSTrack-B256 61.9 60.9 61.4 81.2 88.3 72.8 53.1 66.5 62.5 84.8
UnTrack [17] CVPR’24 OSTrack-B256 61.1 60.8 61.0 82.0 86.9 72.1 51.3 63.7 62.5 84.2
Ours - OSTrack-B256 64.7 65.4 65.1 82.1 88.8 74.1 56.3 70.1 64.4 87.2
ViPT [15] CVPR’23 DropTrack-B384 62.6 61.9 62.2 81.7 87.3 72.3 52.5 65.6 63.4 84.7
SDSTrack [16] CVPR’24 DropTrack-B384 63.3 62.2 62.7 81.4 88.6 73.0 55.1 69.0 65.0 87.1
UnTrack [17] CVPR’24 DropTrack-B384 62.4 61.7 62.0 82.1 87.1 72.2 52.4 65.9 64.1 85.2
Ours - DropTrack-B384 67.2 67.1 67.1 82.9 89.2 74.5 58.8 73.6 66.7 89.1
MamTrack [11] CVPR’25 HiViT-B384 - - - - - - 54.2 67.4 62.4 84.4
SUTrack [8] AAAI’25 SUTrack-B384 58.8 58.6 58.7 83.0 90.6 75.4 60.8 75.6 69.2 92.1
Ours - SUTrack-B384 65.1 65.2 65.2 84.1 92.6 77.7 62.9 77.8 70.3 93.3
TABLE III: Quantitative comparison with OneTracker.
Method Reference Base Model VisEvent DepthTrack VOT RGBD2022 LasHeR RGBT234
SR PR Pr Re F-score Acc Rob EAO SR PR MSR MPR
OneTracker [1] CVPR’24 OSTrack-B384 60.8 76.7 60.7 60.4 60.9 81.9 87.2 72.7 53.8 67.2 64.2 85.7
Ours - OSTrack-B384 62.6 79.4 65.3 66.1 65.6 82.4 88.9 74.3 57.0 71.2 65.4 87.9
IV-B Comparison with State-of-the-Art Methods
Extensive comparative analyses are presented in Tab. I, Tab. II and Tab. III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies. The corresponding precision and success plots are illustrated in Fig. 6, further substantiating the quantitative results. Evidently, we can observe that both the RGB-only and the cross-modal trackers are becoming increasingly profitable with pre-trained models. In particular, cross-modal approaches exhibit substantial performance gains, highlighting the complementarity between RGB and auxiliary data under complex conditions. Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
Results on RGB-Event Tracking. As shown in Tab. I and Tab. III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively. In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256. The full fine-tuning approaches (e.g., CEUTrack, MamTrack) yield limited improvements, while parameter-efficient fine-tuning paradigms (e.g., ViPT, SDSTrack) face performance bottlenecks. This stems from a misfit that impedes cross-modal transfer, emphasizing the effectiveness of our regularized tuning.
Results on RGB-Depth Tracking. As depicted in Tab. II and Tab. III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score. Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score. Similarly, when built on the DropTrack-B384 with richer pre-trained knowledge, our method demonstrates superior performance gains, $3.9\%$ in Pr, $4.9\%$ in Re, and $4.4\%$ in F-score. Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines. This clearly demonstrates the strong generalization and transferability of our approach in cross-modal tracking scenarios.
Results on RGB-Thermal Tracking. As listed in Tab. II and Tab. III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256. Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR. Remarkably, with SUTrack as our pre-trained baseline, our method sets a new state-of-the-art, reaching $77.8\%$ in PR and $62.9\%$ in SR. Beyond the established benchmarks, our method also excels on the unseen RGBT234 dataset, attaining $93.3\%$ MPR and $70.3\%$ MSR. These results not only underscore the superior performance of our method but also validate its exceptional cross-dataset generalization capabilities.
Attribute Analysis. To comprehensively analyze our method, we present a detailed per-attribute comparison in Fig. 7. Our approach consistently achieves the best performance across nearly all challenging attributes while significantly leading. Specifically, for motion-related sequences from the VisEvent and CoeSot datasets, our method delivers the best results, highlighting strong robustness against motion-induced degradation. Particularly, it yields precision gains of $5.6\%$ under Motion Blur, $3.0\%$ under Camera Motion. On the LasHeR, our regularization strategy yields notable improvements under extreme lighting conditions: $13.4\%$ under Illumination Variation, $6.2\%$ under Low Illumination, and $5.3\%$ under Over Exposure. Moreover, our method also maintains superior performance across other challenging attributes such as Out-of-View and Frame Lost.
IV-C Ablation Study and Analysis
We conducted ablative analyses to verify the effectiveness and characteristics of sensitivity-aware regularization tuning. In the ablation section, “full fine-tuning” refers to training the entire backbone, excluding the box head (as Section III-A). All methods use the pre-trained OSTrack-B256 weights unless otherwise specified.
Effectiveness of Proposed Sensitivities. We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Tab. IV. A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability. Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior sensitivity regularization scheme significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness. Similarly, the comparison between (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements. Notably, the simultaneous application of both prior and transfer sensitivity yields significantly greater improvements than either method alone, proving their complementary nature. While the transfer sensitivity penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
TABLE IV: Ablative study results of the proposed key components. “FFT” refers to full fine-tuning of the backbone; (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
Exp. FFT Prior Sensitivity Transfer Sensitivity FE108 DepthTrack LasHeR
SR PR Pr Re F-score SR PR
(a) 48.8 75.2 38.2 36.0 37.1 36.7 43.5
(b) $\checkmark$ 65.2 93.5 61.6 61.4 61.5 53.2 65.8
(c) $\checkmark$ $\checkmark$ 66.8 95.4 64.1 64.5 64.3 55.6 69.0
(d) $\checkmark$ $\checkmark$ 66.6 94.9 63.9 64.2 64.1 55.1 68.4
(e) $\checkmark$ $\checkmark$ $\checkmark$ 67.4 96.5 64.7 65.4 65.1 56.3 70.1
Eigenvalue Budget (Top-$K$) for Prior Sensitivity. In this section, we examine the effect of the low-dimension eigenvalue count $K$ in Eq. (8). For each operation, we record up to the 20 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them. As shown in Tab. V, decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best configuration), and even $K=1$ still yields a clear improvement over the baseline. This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes. Under our low-dimension approximation, the central challenge lies in accurately identifying the operations most sensitive to the generalization gap.
TABLE V: Ablation analysis of low-dimension eigenvalue count $K$ on prior sensitivity estimation.
Exp. FE108 DepthTrack LasHeR
SR PR Pr Re F-score SR PR
KK=1 66.5 95.0 63.9 64.3 64.1 55.2 68.7
KK=2 66.6 95.2 64.0 64.4 64.2 55.4 68.8
KK=5 66.8 95.3 64.1 64.4 64.2 55.5 69.0
KK=10 66.8 95.4 64.1 64.5 64.3 55.6 69.0
KK=20 66.4 95.2 64.1 64.4 64.2 55.4 68.9
Settings of Sensitivity Harmony Coefficient $\kappa$. In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq. (14), with the detailed results presented in Tab. VI. By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9). This finding suggests that while both prior sensitivity and transfer sensitivity contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
TABLE VI: Ablation analysis of sensitivity harmony coefficient on the FE108 dataset.
$\kappa$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
SR 66.5 66.7 66.9 66.9 67.2 67.4 67.1 66.8 66.9
PR 94.7 94.8 95.1 95.3 96.0 96.5 96.2 95.8 95.5
Impact of Box Head Tuning. In our default setting, “full fine-tuning” updates the backbone while keeping the box head frozen. To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well. As shown in Tab. VII, unfreezing the box head further exacerbates over-fitting. While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge encoded during pre-training.
TABLE VII: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the sensitivity-aware regularization tuning method.
Exp. Box FE108 DepthTrack LasHeR
SR PR Pr Re F-score SR PR
FFT w/ 63.8 91.0 59.4 58.7 59.0 51.5 64.4
FFT w/o 65.2 93.5 61.6 61.4 61.5 53.2 65.8
FFT+S-Reg w/ 66.2 95.3 62.6 63.4 63.0 54.4 67.7
FFT+S-Reg w/o 67.4 96.5 64.7 65.4 65.1 56.3 70.1
Compatibility with PEFT Methods. Existing PEFT methods [15, 16, 17] typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization. To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied. As shown in Tab. VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR. Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR. This demonstrates that overly rigid constraints on pre-trained models limit their transfer potential. However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
TABLE VIII: Compatibility study results of the proposed regularization tuning method on existing PEFT methods. “FFT” denotes full fine-tuning of the backbone, while “S-Reg” represents the sensitivity-aware regularization tuning scheme.
Exp. VisEvent DepthTrack LasHeR
SR PR Pr Re F-score SR PR
ViPT 59.2 75.8 59.2 59.6 59.4 52.5 65.1
ViPT+FFT 57.5 74.0 58.2 57.4 57.8 50.9 63.8
ViPT+FFT+S-Reg 61.3 77.8 61.9 61.4 61.6 54.9 68.4
UnTrack 58.9 75.6 61.1 60.8 61.0 51.3 63.7
UnTrack+FFT 56.6 73.1 58.2 56.8 57.5 48.1 60.2
UnTrack+FFT+S-Reg 60.2 76.9 61.9 62.5 62.2 54.0 67.4
SDSTrack 59.7 76.7 61.9 60.9 61.4 53.1 66.5
SDSTrack+FFT 56.0 72.0 57.8 56.4 57.1 50.7 63.8
SDSTrack+FFT+S-Reg 57.8 74.8 59.5 58.3 58.9 52.5 65.8
Ours 61.8 78.6 64.7 65.4 65.1 56.3 70.1
Compatibility with Single-modality Methods. This work aims to mitigate the misfitting issue when adapting the foundation trackers to downstream tasks. A central question is how the proposed regularization methods perform on single-modality data. To investigate this, we conduct ablation studies highlighting their impact on different modalities. As shown in Tab. IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques. For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset. Despite substantial distribution differences, our method significantly and consistently enhances the adaptability of auxiliary modalities across various tasks. These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
TABLE IX: Compatibility study results of the proposed regularization tuning method on single-modal data. “S-Reg” represents the regularization tuning method.
Exp. CoeSot DepthTrack LasHeR
SR PR Pr Re F-score SR PR
RGB 64.3 76.3 53.9 53.0 53.4 47.2 58.6
RGB+S-Reg 68.0 80.0 58.8 58.6 58.7 50.3 62.6
Auxiliary 57.5 69.8 49.0 47.3 48.1 42.7 53.7
Auxiliary+S-Reg 60.5 73.7 52.9 51.3 52.1 45.9 57.8
Figure 8: Operation-wise weight distances $\frac{\left\|\theta_{t}-\theta_{0}\right\|_{2}}{\left\|\theta_{0}\right\|_{2}}$ between the tuned model and the pre-trained model after training, including results from vanilla full fine-tuning (upper) and our regularized tuning (bottom). The auxiliary branches are derived from the VisEvent, DepthTrack, and LasHeR datasets, respectively.
Impact of Event Representations. In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer. In this experiment, we keep the same training setup, varying only the input event representations. As shown in Tab. X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps. In contrast, count-based representations (Event Count, Event Frame) offer more robust performance. The time-interpolation-based Event Volume further improves results by utilizing time-weighted and multi-channel methods to preserve spatio-temporal information. Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
TABLE X: Ablation results of different event representations.
Exp. FE108 VisEvent CoeSot
SR PR SR PR SR PR
Event Frame [64] 66.8 95.5 61.2 77.6 68.8 82.0
Event Count [65] 66.3 95.1 59.8 76.9 67.9 81.2
Time Surface [66] 66.1 94.9 60.3 77.1 67.4 80.6
TSLTD [67] 66.5 95.2 61.0 77.4 67.9 81.3
Event Volume [68] 67.3 96.1 61.4 77.9 69.1 82.4
Color-Polar Event Image (Ours) 67.5 96.5 61.8 78.6 69.6 83.0
Effectiveness of Smaller Learning Rates. In this paper, we exploit the sensitivity-regularized gradients to bolster the transfer process. A straightforward strategy might involve using a smaller learning rate. To investigate this, we evaluate the effect of reduced learning rates. As shown in Tab. XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR. Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation. These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
TABLE XI: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
Exp. SR PR
$\alpha=10^{-4}$ 53.2 65.8
$\alpha=10^{-5}$ 52.9 66.2
$\alpha=10^{-6}$ 52.5 65.4
Ours 56.3 70.1
Observations on Weight Variations. To intuitively demonstrate the effectiveness of the proposed sensitivity-aware regularization tuning, we visualize parameter dynamics during training from both spatial and temporal perspectives, as depicted in Fig. 8 and Fig. 9. Overall, our regularization encourages smaller and more evenly distributed weight deviations, effectively mitigating over-fitting and facilitating stable adaptation. As shown in Fig. 8, for parameters with high prior sensitivity, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates. As further demonstrated in Fig. 9, our method facilitates faster convergence and achieves a substantial reduction in sensitivity-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation. Moreover, the vanilla fine-tuning causes parameter updates to concentrate excessively in localized regions, hindering the global adaptability. After applying the sensitivity-aware penalties, the parameter updates become more balanced and distributed, effectively decentralizing the adaptation burden from a few parameters to a broader set. This redistribution helps mitigate adaptation risk and promotes more stable fine-tuning.
Figure 9: Sensitivity-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
TABLE XII: Computational complexity and speed analysis on the LasHeR dataset.
Method Base Model Param (M) Flops (G) FPS SR PR
OSTrack [5] OSTrack 92.1 58.1 98.4 47.2 58.5
ProTrack [41] OSTrack 92.7 58.4 92.3 42.0 53.8
ViPT [15] OSTrack 92.9 59.9 88.6 52.5 65.1
SDSTrack [16] OSTrack 102.1 108.7 44.6 53.1 66.5
UnTrack [17] OSTrack 98.7 62.6 75.6 51.3 63.7
ViPT+Ours OSTrack 92.9 59.9 88.6 54.9 68.4
UnTrack+Ours OSTrack 98.7 62.6 75.6 54.0 67.4
Ours OSTrack 202.0 149.0 49.0 56.3 70.1
DropTrack [40] DropTrack 92.1 130.6 57.6 47.7 59.2
ViPT [15] DropTrack 92.9 131.9 48.8 52.5 65.6
SDSTrack [16] DropTrack 102.1 244.2 26.8 55.1 69.0
UnTrack [17] DropTrack 98.7 138.2 42.3 52.4 65.9
ViPT+Ours DropTrack 92.9 131.9 48.8 57.1 70.5
UnTrack+Ours DropTrack 98.7 138.2 42.3 56.3 69.4
Ours DropTrack 202.0 335.3 29.2 58.8 73.6
SUTrack [8] SUTrack 65.29 106.9 37.6 60.8 75.6
Ours SUTrack 141.1 243.7 24.8 62.9 77.8
Figure 10: Visual comparisons of the tracking performance of different methods on the (Left) RGB-Event, (Middle) RGB-Depth and (Right) RGB-Thermal datasets.
IV-D Computational Efficiency
In addition to tracking accuracy, we evaluate the computational efficiency of our sensitivity-aware regularization tuning by analyzing both training and inference overhead.
Efficiency of Regularized Fine-tuning. We compare the training efficiency of our method against vanilla fine-tuning. Since prior sensitivity estimation is conducted as a one-time preprocessing step before fine-tuning, its runtime is excluded from the training time. Our method introduces parameter-wise transfer sensitivity quantification with marginal computational overhead, as this sensitivity is derived from the off-the-shelf gradients. While vanilla fine-tuning operates at 37.5 ms per iteration, our method runs at 39.0 ms per iteration, incurring only a modest $4.0\%$ increase in training time.
Inference Cost and Speed. We compare the computational complexity and runtime efficiency of existing methods, focusing on models built upon OSTrack, DropTrack and SUTrack. Note that the proposed regularization techniques are applied only during training, imposing no additional computational burden during testing. All experiments are conducted on the same system, which includes an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and one NVIDIA 3090Ti GPU. As reported in Tab. XII, our method supports real-time tracking at 24.8 frames per second, while delivering top-tier performance. When integrated into ViPT, our approach (“ViPT+Ours”) yields both superior speed and performance. Under the DropTrack base model, it still operates at 48.8 frames per second, maintaining a respectable speed given its focus on accuracy.
IV-E Visual Comparison
Some representative visualization results are illustrated in Fig. 10. For clarity, all visual instances are explicitly labeled by the predicted bounding box. Horizontally, each instance is represented by a pair of pictures - the right image shows the main view, while the left displays the auxiliary modality. Our method demonstrates a marked improvement over state-of-the-art multi-modal methods, especially in handling complex scenarios where cross-modal integration plays a crucial role. It excels in tracking in conditions like motion blur, occlusions, and low illumination, where other methods struggle. These results demonstrate that our method improves tracking accuracy and provides a more stable, efficient way to integrate auxiliary data.
V Conclusion
This study revisits the critical misfitting issues encountered when adapting pre-trained RGB trackers to multi-modal tracking tasks. By introducing two complementary sensitivities that capture the dynamic shift in parameter importance as models transition from pre-trained to multi-modal contexts, our regularized tuning method strategically calibrates gradient updates. Extensive experiments demonstrate the effectiveness of our method, surpassing current state-of-the-art techniques across various multi-modal tracking scenarios, with notable performance improvements resulting from the incorporation of regularization. A key insight from our work is the importance of sensitivity-aware fine-tuning, which offers superior generalization and adaptability compared to traditional fine-tuning methods. Ultimately, our findings highlight the need for more nuanced cross-domain generalization strategies, as overly rigid or flexible fine-tuning can hinder pre-trained model transferability. We believe these insights will pave the way for further advancements in multi-modal transfer learning, particularly in the context of scene perception tasks.
Proof of Proposition 1.
  1. 1.
    Bounded FIM Error. Each group approximation $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0}^{j})}$ with bounded error. Due to symmetry and orthogonal eigenvectors, the Frobenius norm of the error is derived from the difference in their eigenvalues:
    $\|\mathcal{F}^{(\theta_{0}^{j})}-\tilde{\mathcal{F}}^{(\theta_{0}^{j})}\|_{F}=\sqrt{\sum_{i=1}^{|\theta^{j}|}(\lambda_{i}^{j}-\gamma^{j})^{2}},$
    Since $\gamma^{j}$ lies between the largest and smallest eigenvalues of $\mathcal{F}^{(\theta_{0}^{j})}$, i.e., $\lambda_{|\theta^{j}|}^{j}<\gamma^{j}<\lambda_{1}^{j}$, we can deduce that:
    $\|\mathcal{F}^{(\theta_{0}^{j})}-\tilde{\mathcal{F}}^{(\theta_{0}^{j})}\|_{F}\leq\sqrt{\sum_{i=K+1}^{|\theta^{j}|}{(\lambda_{i}^{j})}^{2}},$
    Summing over all groups, the total discarded eigenvalue mass is:
    $\|\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})}\|_{F}\leq\sqrt{\sum_{j=1}^{N}\sum_{i=K+1}^{|\theta^{j}|}{(\lambda_{i}^{j})}^{2}},$
    In particular, under the common assumption that the top-$K$ eigenvalues in each group capture the most significant curvature, where $\mathrm{Tr}(\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})})\ll\mathrm{Tr}(\mathcal{F}^{(\theta_{0})})$, the residual eigenvalue mass is small.
  2. 2.
    Bounded Generalization Gap Error. The approximate FIM $\tilde{\mathcal{F}}^{(\theta_{0})}$ remains close to $\mathcal{F}^{(\theta_{0})}$ as a metric on parameter space. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded in terms of the residual generalization gap error. In particular, one has following formula by definition of the spectral norm:
    $|\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})-\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})|\leq\frac{1}{2}\|\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})}\|_{2}\,\|\Delta\theta\|^{2},$
    Since $\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})}$ is group-diagonal (with each block $\mathcal{F}^{(\theta_{0}^{j})}-\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$) and symmetric, its spectral norm is:
    $\|\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})}\|_{2}=\underset{1\leq j\leq N}{\operatorname{\max}}\|\mathcal{F}^{(\theta_{0}^{j})}-\tilde{\mathcal{F}}^{(\theta_{0}^{j})}\|_{2},$
    Using the bound above, we obtain
    $|\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})-\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})|\leq\frac{1}{2}\|\Delta\theta\|^{2}(\underset{1\leq j\leq N}{\operatorname{\max}}\lambda_{1}^{j}),$
    Specifically, due to the general principle that the spectral norm of a symmetric matrix is always smaller than its Frobenius norm, we can also derive:
    $|\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})-\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})|\leq\frac{1}{2}\|\Delta\theta\|^{2}\sqrt{\sum_{i=K+1}^{|\theta^{j}|}{(\lambda_{i}^{j})}^{2}},$
    In practice, by retaining the top-$K$ eigenvalues in each FIM block (which capture the majority of the FIM energy) and replacing the remaining eigenvalues with an isotropic average, one obtains a low-rank Fisher matrix approximation that preserves the important Riemannian geometry of the parameter space while bounding the distortion in any distance or generalization metric by the residual eigenvalue mass.

### v2

@@
-Optimizing Multi-Modal Trackers via Sensitivity-aware Regularized Tuning
+Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning
 Abstract
-This paper tackles the critical challenge of optimizing multi-modal trackers by effectively adapting the pre-trained models for RGB data.
+This paper tackles the critical challenge of optimizing multi-modality trackers by effectively adapting pre-trained models for RGB data.
 Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off.
-To mitigate this dilemma, we propose a novel sensitivity-aware regularized tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities.
-Through a comprehensive investigation from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are primary drivers of this issue.
-Specifically, we first analyze the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization.
-Then, we further explore transfer sensitivities during the tuning phase, emphasizing adaptability and stability.
-By incorporating these sensitivities as regularization terms, our method significantly enhances the transferability across modalities.
-Extensive experiments showcase the superior performance of the proposed method, surpassing current state-of-the-art techniques across various multi-modal tracking.
+To mitigate this dilemma, we propose a novel sensitivity-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities.
+Through a comprehensive investigation of the transition from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are the primary drivers of this issue.
+Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization.
+Subsequently, we characterize transfer sensitivities during the tuning phase, emphasizing adaptability and stability.
+By incorporating these sensitivities as unified regularization terms, our method significantly enhances the transferability across modalities.
+Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modality tracking benchmarks.
 The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
-Index Terms:
-Multi-modal Tracking, Cross-modal Transfer, Parameter Sensitivity, Regularized Tuning.
-Figure 1: Training and testing performance on LasHeR, comparing our method with existing fine-tuning approaches.
-Our method effectively mitigates the misfitting issue and enhances both generalization and stability of the multi-modal tracker.
+Figure 1: Optimization trajectory analysis on LasHeR.
+This plot contrasts the training and testing dynamics of different tuning paradigms.
+As visualized by the colorful arrows, our method effectively mitigates the misfitting issue and enhances multi-modality trackers with superior generalization and stability.
 I Introduction
@@
 Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions.
-Therefore, multi-modal tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest.
+Therefore, multi-modality tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest.
 With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [58, 29, 10, 9].
 There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
-Some existing approaches have explored the full fine-tuning (FFT) paradigm [9, 10, 11], where these trackers are initialized with pre-trained weights and subsequently tuned for task-specific objectives.
-Nevertheless, due to the significant gap and limited scale of auxiliary modalities, they are intractable to retain the pre-trained knowledge during transfer and tend to induce catastrophic forgetting and overfitting (i.e., widened train-test gap).
+Some existing approaches have explored the full fine-tuning (FFT) paradigm [45, 73, 41], where trackers are initialized with pre-trained weights and subsequently optimized for task-specific objectives.
+Nevertheless, due to the significant domain gaps and limited scale of auxiliary modalities, these methods struggle to preserve the pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance).
 In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [18, 21, 7].
-The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modal-specific ones to retain pre-trained knowledge.
+The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modality-specific ones to retain prior knowledge.
 Several methods fall under this umbrella [70, 17, 51], including prompt tuning, visual adapter, etc.
-While effective, PEFT methods impose excessive constraints on the primary model weights, resulting in underfitting (i.e., restricted train upper bound) for handling the vast distribution drift.
-In summary, existing tuning paradigms fluctuate between excessive flexibility and overly rigid constraints, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and adaptation (please refer to Fig.
+Although effective, PEFT-based methods impose rigid constraints on the primary model weights, resulting in underfitting (i.e., a restricted upper bound on training performance) when handling the significant distribution drifts.
+In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as Fig.
 1).
-In this work, we endeavor to mitigate the misfitting issue in cross-modal tracker adaptation by strategically regularizing the learning process.
-By examining parameter responses spanning from the pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability.
-To mitigate this, we propose a sensitivity-aware regularized tuning framework that delicately calibrates the gradient updates to boost model transfer with precision.
-Specifically, we optimize multi-modal trackers from the following perspectives.
+In this work, we endeavor to mitigate the misfitting dilemma in cross-modal tracker adaptation by strategically regularizing the learning process.
+By probing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability.
+To mitigate this, we propose a sensitivity-regularized fine-tuning framework that delicately calibrates the gradient updates to boost model transfer with precision.
+Specifically, we optimize multi-modality trackers from the following perspectives.
   1. (1)
     Formulating Prior Sensitivity.
-We begin by investigating the tangent space of pre-trained parameters, using it as a critical indicator to assess and preserve prior generalizability.
-Moreover, we estimate such prior sensitivity via an eigen-decomposition approximation.
+We commence by investigating the tangent space of pre-trained parameters, utilizing it as a critical indicator to assess and preserve prior generalizability for downstream fine-tuning.
+Furthermore, we estimate this sensitivity via an eigen-decomposition approximation.
   2. (2)
     Modeling Transfer Sensitivity.
-To further elucidate transfer challenges, we explore how sparse gradients exacerbate tuning instability.
-Accordingly, we construct this transfer sensitivity using the off-the-shelf gradient matrix, with the aim of facilitating gradient rebalancing.
+To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate tuning instability.
+Accordingly, we formulate transfer sensitivity using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
   3. (3)
-    Sensitivity-Aware Regularization Tuning.
-By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modal representations through finely modulated, sensitivity-driven updates.
-This regularization tuning facilitates seamless transfer to various multi-modal tracking tasks, which can continuously enhance the model in the training phase.
-Our method strategically guides the cross-modal fine-tuning process to optimize downstream tracking tasks.
-Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modal tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models.
+    Sensitivity Regularized Tuning.
+By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, sensitivity-driven updates.
+This mechanism facilitates seamless transfer across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
+Our method strategically guides the cross-domain fine-tuning process to optimize downstream tracking tasks.
+Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modality tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models.
 Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
 In summary, the main contributions of this paper are:
-  • we revisit the misfitting issue of multi-modal tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
-  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of multi-modal trackers;
-  • we conduct comprehensive experiments covering three multi-modal tracking tasks and seven datasets using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels;
+  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
+  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
+  • we conduct comprehensive experiments covering three multi-modality tracking tasks and seven benchmarks using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels.
 The remainder of the paper is organized as follows.
-Section II reviews existing works on multi-modal trackers and cross-modal transfer learning methods.
-In Section III, we explicitly formulate the prior and transfer sensitivity from both the pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-aware regularized tuning.
-In Section IV, we present experiments on seven multi-modal tracking benchmarks to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs.
+Section II reviews existing literature on multi-modal trackers and cross-modal transfer learning methods.
+In Section III, we rigorously formulate prior and transfer sensitivities from both pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-regularized tuning.
+Section IV presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs.
 Finally, Section V concludes the paper.
 II Related Work
-II-A Multi-Modal Tracking
+II-A Multi-Modal Object Tracking
 Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [48, 2, 28, 68].
-In recent years, due to the unreliability of RGB data in challenging scenarios, increasing studies are expected to integrate auxiliary modalities (e.g., event, depth, thermal data) to enhance tracking performance.
-These complementary signals compensate for the inherent limitations of RGB imagery [22].
+However, the unreliability of RGB-only paradigms to adverse conditions has catalyzed research into multi-modal tracking, where auxiliary signals (e.g., event, depth, thermal data) are integrated to remedy the intrinsic deficiencies of visible imagery [66].
 For instance, event cameras provide robust dynamic scene information under extreme motion or lighting.
-Leveraging this, [23, 24, 25, 26, 27, 28] combine RGB and event streams to predict objects in high-speed and low-dynamic scenarios.
-Meanwhile, depth data offers spatial cues that are particularly useful for handling occlusion and background clutter.
-[29, 30, 31, 32] exploit this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes.
-Similarly, thermal infrared imaging captures heat signatures independent of visible light, proves effective in low-illumination.
-[33, 34, 35, 36] demonstrate that fusing thermal and RGB data can yield more reliable appearance representations.
-In summary, these methods emphasize effective multi-modal interaction and fusion [37, 9, 38, 39].
-With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [5, 40, 8] have achieved remarkable generalization across diverse scenarios.
+Leveraging this, [72, 11, 46, 63, 43, 44] combined RGB and event streams to predict objects in high-speed and low-dynamic scenarios.
+Meanwhile, depth cues offer geometric priors that are particularly useful for handling occlusion and background clutter.
+[32, 36, 56, 31] exploited this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes.
+Similarly, thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination.
+[62, 19, 27, 52] demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
+In summary, contemporary methods emphasized effective multi-modal interaction and fusion [42, 45, 59, 65].
+With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [58, 49, 9] have demonstrated remarkable generalization across diverse scenarios.
 These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers.
-Multi-modal tracking is increasingly driven by reusing RGB pre-trained semantic priors, particularly when annotated multi-modal data are scarce or noisy.
-In this work, we target this challenge by focusing on optimizing the adaptation of pre-trained trackers for efficient cross-modal transfer.
+Consequently, multi-modal tracking is increasingly driven by reusing RGB-induced semantic priors, particularly when annotated multi-modal data are scarce or noisy.
+In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modality transfer.
 II-B Cross-modal Transfer Learning
 To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged.
-Some works follow the full fine-tuning (FFT) paradigm [37, 9, 10, 11], where the pre-trained models serve as initializations and are entirely re-trained on the target tasks.
+Some works adhere to the full fine-tuning (FFT) paradigm [42, 45, 73, 41], where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks.
 These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model.
-Representatively, SUTrack [8] employs a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization.
-While often effective, one primary drawback may be innate to this paradigm: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation often leads to significant catastrophic forgetting or overfitting.
-To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [13, 14, 12], some parameter-efficient fine-tuning (PEFT) methods have been proposed.
+Representatively, SUTrack [9] employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization.
+While effective, one primary dilemma may be innate to FFT: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation.
+This imbalance frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data.
+To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [21, 7, 18], the field has pivoted toward parameter-efficient fine-tuning (PEFT).
 The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task.
 By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting.
-For instance, ProTrack [41] and ViPT [15] modulate RGB features by introducing trainable auxiliary tokens into attention layers.
-Similarly, methods like BAT [42] and SDSTrack [16] insert lightweight adapter modules between attention layers for cross-modal shift compensation.
-More recently, unified trackers like UnTrack [17] and OneTracker [1] employ LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs.
-Despite their effectiveness, excessive constraints on pre-trained weights can incur insufficient adaptation.
-Furthermore, the added modules tend to over-fit since the additional parameters are optimized from scratch when tailored to modality-specific tasks.
-Consequently, how to efficiently fine-tune RGB-based pre-trained models for multi-modal tracking remains a significant challenge.
-Figure 2: Model architecture for our multi-modality tracker.
+For example, ProTrack [57] and ViPT [70] modulated RGB features by introducing trainable auxiliary tokens into attention layers.
+Similarly, methods like BAT [5] and SDSTrack [17] inserted lightweight adapter modules between attention layers for cross-modal shift compensation.
+More recently, unified trackers like UnTrack [51] and OneTracker [16] employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs.
+Despite their effectiveness, PEFT methods suffer from intrinsic limitations.
+First, the rigid constraints can lead to underfitting, limiting the model’s capacity to handle vast distribution drifts.
+Conversely, the randomly initialized add-on modules lack alignment with the pre-trained manifold, creating an optimization gap that predisposes the model to overfitting.
+Consequently, achieving robust and reliable fine-tuning for multi-modal tracking remains an elusive goal.
+Remark.
+While recent works have explored sensitivity-aware mechanisms in efficient fine-tuning (SPT) and continual learning (CL), our SRFT introduces a fundamentally distinct paradigm in objectives and mechanisms.
+Specifically, methods like SPT [15] and CDRA-SPT [8] utilized sensitivity to prioritize structural adaptation, such as parameter masking or dynamic rank allocation for parameter efficiency.
+Conversely, MACL [47] employed sensitivity to prevent forgetting in sequential tasks by minimizing loss variance.
+Crucially, these approaches typically operate with an exclusive focus on the target or current data.
+Diverging from this, SRFT establishes a gradient-regularized paradigm.
+We formulate sensitivity to balance the preservation of pre-trained knowledge with sparse transfer gradient, via a hybrid sensitivity from pre-trained to multi-modal contexts.
+This distinction renders our method orthogonal to existing FFT, PEFT, and SPT frameworks.
+Figure 2: Network architecture of the multi-modality trackers.
+All modules are initialized with the weights of a pre-trained RGB tracker.
+We only fine-tune backbones and fusion blocks with sensitivity-aware regularization terms.
 III Proposed Method
 Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers.
-To unlock the full potential of pre-trained trackers, we revisit the core principles of cross-modal transfer process.
+To unlock the full potential of pre-trained trackers, we revisit the core principles governing the cross-modal transfer process.
 In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning.
-Next, in Section III-B, we identify pronounced prior and transfer sensitivities that contribute to catastrophic forgetting and training instability, and further quantify these critical sensitivities.
-Finally, in Section III-C, we incorporate these sensitivities into the tuning phase, enabling adaptive and dynamic modulation of parameter updates.
+Next, Section III-B identifies and quantifies pronounced prior and transfer sensitivities that characterize pre-trained knowledge and cross-domain adaptation.
+Finally, in Section III-C, we incorporate these sensitivities into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
 III-A Preliminaries
-Network Architecture.
+Network Architecture of Multi-modality Trackers.
 Fig.
-2 depicts the overall architecture of our multi-modal tracker.
-The RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens.
+2 depicts the overall architecture of our multi-modal tracker, in which all modules are initialized with the weights of a pre-trained RGB tracker.
+For a multi-modal tracker, the RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens.
 Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction.
 Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens.
+Finally, the fused features are fed into the box head to estimate the object state.
 To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
-Cross-Domain Transfer.
-We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameter $\theta$.
+Learning Process of Cross-Domain Transfer.
+We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameters $\theta$.
 It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box.
@@
 Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
-$\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\theta^{(0)}=\theta_{0},$ (1)
+$\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\quad\theta^{(0)}=\theta_{0},$ (1)
 However, directly optimizing $\theta_{t}$ as per Eq.
-(1) leads to severe catastrophic forgetting and unstable tuning, as evidenced by the pronounced over-fitting and oscillations in training dynamics observed in Fig.
+(1) leads to severe overfitting and unstable tuning, as evidenced by the pronounced train-test performance gap and oscillations in training dynamics observed in Fig.
 1.
-This motivates a more principled transfer learning approach described next.
-Figure 3: Loss-parameter manifold schematic for different fine-tuning paradigms.
-The “FFT” fine-tunes all pre-trained weights without regularization, leading to significant catastrophic forgetting on pre-trained dataset.
-Although “PEFT”-based methods only tune the additional parameters to maintain the pre-traiend knowledge, but their coarse modeling limits performance on new domain.
-In contrast, our “SRFT” method carefully optimize the model within an approximated tangent space of the pre-trained manifold, indicating a better plasticity-stability trade-off between pre-trained knowledge and adaptation.
-III-B Modeling Sensitivity for Multi-Modal Tracking
+This motivates a more principled transfer learning approach, as described next.
+Figure 3: Loss-parameter manifold schematic across different fine-tuning paradigms.
+FFT updates all weights without constraints, risking severe forgetting of pre-trained knowledge (deteriorating pre-trained loss, L1).
+While PEFT restricts updates to additional parameters, this limits performance on new domains (stalled transferred loss, L2).
+In contrast, our SRFT performs optimization within an approximated pre-trained tangent space, indicating a better plasticity-stability trade-off.
+III-B Modeling Sensitivity for Multi-Modality Tracking
 Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) [34].
@@
 3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$.
-Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$.
-Based on the aforementioned manifold hypothesis, to avoid the losing of generalization ability for fine-tuned models on downstream domains, the critical thing lies on optimize the model on local tangent space, shown as the red arrow of the Fig.
+Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ [13, 40].
+Based on the aforementioned manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig.
 3.
@@
 $\approx\frac{1}{2}\left(\theta-\theta_{0}\right)^{T}\mathcal{F}^{\left(\theta_{0}\right)}\left(\theta-\theta_{0}\right)=\frac{1}{2}\left\|\theta-\theta_{0}\right\|_{\mathcal{F}^{\left(\theta_{0}\right)}}^{2},$
-This generalization gap is measured as the weight distance within the Riemannian manifold defined by the FIM.
-Through geometric perspective, the FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [45, 46].
-Therefore, the FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
-Figure 4: The operation-wise eigen-decomposition FIM map (i.e., prior sensitivity) of the pre-trained OSTrack.
+This generalization gap is measured as the weight distance within the Riemannian manifold defined by FIM.
+Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [30, 50].
+Therefore, FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
+Figure 4: Operation-wise prior sensitivity (i.e., eigen-decomposition FIM map) of the pre-trained OSTrack.
 The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
-Eigen-decomposition FIM Approximation.
-Directly computing the FIM for a large model is computationally intractable due to its $\mathcal{O}\left(|\theta|^{2}\right)$ complexity.
-Instead, we propose constructing an eigen-decomposition approximation of the FIM without explicitly forming it in full.
-Focusing on leading eigenvalues captures key tangent geometrics of affecting generalization while reducing complexity.
+Eigen-decomposition Approximation of FIM.
+Directly computing FIM for a large model is computationally intractable due to its $\mathcal{O}\left(|\theta|^{2}\right)$ complexity.
+Instead, we propose to approximate its eigen-decompositionwithout explicitly forming the full matrix.
+Intuitively, this approximation can be likened to analyzing the terrain of the pre-trained loss landscape.
+Eigenvectors associated with large eigenvalues represent “steep cliffs” directions where modifying parameters would drastically increase the pre-trained loss and destroy prior knowledge.
+Conversely, small eigenvalues represent “flat valleys” regions where parameters can be adjusted freely to adapt to new modalities without harming the foundation model.
+By focusing on leading eigenvalues, we efficiently identify the most critical parameter subspaces that must be preserved, while reducing overall complexity.
 Specifically, we partition the model’s parameters into $N$ disjoint operation groups, $\theta=\left\{\theta^{1},\cdots,\theta^{N}\right\}$ (e.g.
@@
 Parameters from different groups thus span orthogonal subspaces with no second-order coupling.
-Empirical studies [47, 48] have shown that the FIM of deep networks are spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs.
+Empirical studies [14, 37] have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs.
 Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
@@
 Proof 1
-See Appendix.
+See Appendix I-A.
 Prior Sensitivity Measure.
 Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [23].
-Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to Rayleigh quotient of FIM.
+Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to the Rayleigh quotient of FIM.
 This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters.
@@
 $\left\|\mathcal{G}\right\|_{2}=\frac{\left\|\mathcal{G}\right\|_{1}}{\rho\sqrt{|\theta|}}.$ (9)
-Figure 5: The instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning.
+Figure 5: Instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning.
 The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
@@
 $S\left(\theta,\delta\mid M\right)=\mathcal{L}\left(\theta\mid M\right)-\mathcal{L}\left(\theta+\delta\mid M\right),$ (10)
-where $\delta$ is a random perturbation centered on the current update step, i.e., $\delta\sim\mathcal{N}\left(\alpha\mathcal{G},\sigma^{2}I\right)$, $\alpha$ is the learning rate and $\sigma$ controls the noise.
+where $\delta$ is a random perturbation centered on the current update step, i.e., $\delta\sim\mathcal{N}\left(\alpha\mathcal{G},\alpha^{2}\sigma^{2}I\right)$, $\alpha$ is the learning rate and $\sigma$ controls the noise.
 This denotes that updates follow the gradient with inherent noise (due to stochastic sampling, batch gap, etc.).
@@
 By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq. (11), it can be shown that this risk is proportional to the $L_{2}$ norm of gradient:
-$\varepsilon_{ada}\approx\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
+$\varepsilon_{ada}\leq\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
+Derivations details of Eq.
+(9)-(12) can be found in Appendix I-B.
 Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk.
@@
 To further investigate and regularize gradients, we extend the sensitivity to a parameter-wise formulation:
-$s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}})^{2},$ (13)
+$s_{n}^{t}=\left(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}}\right)^{2}.$ (13)
 where $s_{n}^{t}$ denotes the transfer sensitivity of the $n^{\text{th}}$ parameter.
 This formulation allows us to examine the granular impact of parameter changes during the transfer process and rebalance the gradient accordingly.
-1Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
-2Output: Optimal multi-modal tracker parameters $\theta_{t}$;
-3Step1: Prior Sensitivity Estimation:
+Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
+Output: Optimal multi-modal tracker parameters $\theta_{t}$;
+Step1: Prior Sensitivity Estimation:
 for batch $(x,y)$ in $D_{0}$ do
@@
 end for
-4 Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
-5Step2: Transfer Sensitivity Estimation and Regularized Tuning:
+ Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
+Step2: Transfer Sensitivity Estimation and Regularized Tuning:
 for $i\in\{1,\ldots,T\}$ do
   Sample $i$-th batch data $M_{i}$ from $D_{t}$;
-6  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
-7  Update transfr sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
-8  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
-9  Parameter update via Eq. (15);
+  Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
+  Update transfer sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
+  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
+  Parameter update via Eq. (15);
 end for
-Algorithm 1 Sensitivity-Aware Regularization Tuning
-III-C Sensitivity-Aware Regularization Tuning
+Algorithm 1 Sensitivity-Regularized Fine-Tuning
+III-C Sensitivity-regularized Tuning
 We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers.
@@
 This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability.
-Formally, at each training step $t$, the combined sensitivity is updated as follows:
-$s_{n}=(\kappa+(1-2\kappa)\frac{t}{T})s_{j}^{p}+(1-\kappa-(1-2\kappa)\frac{t}{T})s_{n}^{t},$ (14)
+Formally, at each training step $i$, the combined sensitivity is updated as follows:
+$s_{n}=\left(\kappa+(1-2\kappa)\frac{i}{T}\right)s_{j}^{p}+\left(1-\kappa-(1-2\kappa)\frac{i}{T}\right)s_{n}^{t},$ (14)
 where $\theta_{n}\in\theta^{j}$, $T$ is the total number of training steps.
@@
 $\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\text{s.t. }\theta^{(0)}=\theta_{0},$ (15)
-where $\alpha$ is the learning rate.
-This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
-The detailed regularization tuning process is given in Algorithm.1.
-We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments.
-Further comprehensive analysis can be found in Tab.
-VI.
+where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
+The detailed regularization tuning process is given in Algorithm 1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table VI.
 III-D Learning Objectives
@@
 where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as [58].
-Figure 6: Visualization of the precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
+Figure 6: Precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
 Figure 7: Precision scores of different attributes on the VisEvent, CoeSot and LasHeR test sets. Zoom in to see details.
 IV Experiments
-IV-A Experimental Settings
+IV-A Experiment Settings
 IV-A1 Benchmark Datasets
@@
 IV-A2 Evaluation Metrics
-Adhering to recognized standards [16, 17, 1, 11], we evaluated tracking performance with following metrics.
+Adhering to recognized standards [17, 51, 16, 41], we evaluated tracking performance with the following metrics.
 For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR).
@@
 Notably, OSTrack and DropTrack are built upon ViT-B/16 [12], while SUTrack adopts the HiViT-B/16 [64] as the backbone.
-Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack and SUTrack ($\times192$$\times384$) for template and search regions, respectively.
-To comprehensively validate the effectiveness of our method, we conduct the following experiments in Tab.
-I and Tab.
-II.
+Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack, and SUTrack ($\times192$$\times384$) for template and search regions, respectively.
+To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table I and Table II.
 First, we construct strong RGB-based single-modal baselines under full fine-tuning.
@@
 Notably, OneTracker [16] and MamTrack [41] adopt non-standard experimental settings and lack open-source implementations.
-OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Tab.
-III.
+OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table III.
 MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
@@
 IV-B Comparison with State-of-the-Art Methods
-Extensive comparative analyses are presented in Tab.
-I, Tab.
-II and Tab.
-III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies.
+Extensive comparative analyses are presented in Table I, Table II and Table III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies.
 The corresponding precision and success plots are illustrated in Fig.
@@
 Results on RGB-Event Tracking.
-As shown in Tab.
-I and Tab.
-III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively.
+As shown in Table I and Table III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively.
 In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256.
@@
 Results on RGB-Depth Tracking.
-As depicted in Tab.
-II and Tab.
-III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score.
+As depicted in Table II and Table III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score.
 Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score.
@@
 Results on RGB-Thermal Tracking.
-As listed in Tab.
-II and Tab.
-III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256.
+As listed in Table II and Table III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256.
 Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR.
@@
 Effectiveness of Proposed Sensitivities.
-We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Tab.
-IV.
+We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Table IV.
 A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability.
@@
 TABLE IV: Ablative study results of the proposed key components.
-“FFT” refers to full fine-tuning of the backbone; (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
-Exp. FFT Prior Sensitivity Transfer Sensitivity FE108 DepthTrack LasHeR
+“FFT” refers to full fine-tuning of the backbone; “PS” represents prior sensitivity; “TS” is transfer sensitivity.
+(a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
+Exp. FFT PS TS FE108 DepthTrack LasHeR
 SR PR Pr Re F-score SR PR
@@
 (e) $\checkmark$ $\checkmark$ $\checkmark$ 67.4 96.5 64.7 65.4 65.1 56.3 70.1
+Figure 8: Operation-wise eigenvalue spectra for measuring prior sensitivity.
+Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale.
+For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
 Eigenvalue Budget (Top-$K$) for Prior Sensitivity.
-In this section, we examine the effect of the low-dimension eigenvalue count $K$ in Eq.
+In this section, we examine the effect of the eigenvalue count $K$ in Eq.
 (8).
-For each operation, we record up to the 20 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them.
-As shown in Tab.
-V, decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best configuration), and even $K=1$ still yields a clear improvement over the baseline.
+For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them.
+As illustrated in Fig.
+8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues.
+Consistent with this observation, Table V shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline.
 This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes.
-Under our low-dimension approximation, the central challenge lies in accurately identifying the operations most sensitive to the generalization gap.
+Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
 TABLE V: Ablation analysis of low-dimension eigenvalue count $K$ on prior sensitivity estimation.
@@
 KK=20 66.4 95.2 64.1 64.4 64.2 55.4 68.9
-Settings of Sensitivity Harmony Coefficient $\kappa$.
+Sensitivity Harmony Coefficient $\kappa$.
 In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq.
-(14), with the detailed results presented in Tab.
-VI.
+(14), with the detailed results presented in Table VI.
 By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9).
@@
 To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well.
-As shown in Tab.
-VII, unfreezing the box head further exacerbates over-fitting.
-While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge encoded during pre-training.
+As shown in Table VII, unfreezing the box head further exacerbates overfitting.
+While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge.
 TABLE VII: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the sensitivity-aware regularization tuning method.
@@
 To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied.
-As shown in Tab.
-VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR.
+As shown in Table VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR.
 Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR.
@@
 To investigate this, we conduct ablation studies highlighting their impact on different modalities.
-As shown in Tab.
-IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques.
+As shown in Table IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques.
 For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset.
@@
 In this experiment, we keep the same training setup, varying only the input event representations.
-As shown in Tab.
-X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps.
+As shown in Table X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps.
 In contrast, count-based representations (Event Count, Event Frame) offer more robust performance.
@@
 To investigate this, we evaluate the effect of reduced learning rates.
-As shown in Tab.
-XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR.
+As shown in Table XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR.
 Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation.
@@
 TABLE XI: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
-Exp. SR PR
-$\alpha=10^{-4}$ 53.2 65.8
-$\alpha=10^{-5}$ 52.9 66.2
-$\alpha=10^{-6}$ 52.5 65.4
-Ours 56.3 70.1
+Exp. $\alpha=10^{-4}$ $\alpha=10^{-5}$ $\alpha=10^{-6}$ Ours
+SR 53.2 52.9 52.5 56.3
+PR 65.8 66.2 65.4 70.1
 Observations on Weight Variations.
@@
 9 and Fig.
-9.
+10.
 Overall, our regularization encourages smaller and more evenly distributed weight deviations, effectively mitigating over-fitting and facilitating stable adaptation.
 As shown in Fig.
-8, for parameters with high prior sensitivity, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates.
+9, for parameters with high prior sensitivity, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates.
 As further demonstrated in Fig.
-9, our method facilitates faster convergence and achieves a substantial reduction in sensitivity-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation.
+10, our method facilitates faster convergence and achieves a substantial reduction in sensitivity-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation.
 Moreover, the vanilla fine-tuning causes parameter updates to concentrate excessively in localized regions, hindering the global adaptability.
@@
 Figure 10: Sensitivity-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
-TABLE XII: Computational complexity and speed analysis on the LasHeR dataset.
+TABLE XII: Computational cost of Prior Sensitivity Estimation on OSTrack-B256.
+Metric Computational Cost
+Operation-wise Model-wise Total Process
+FLOPs 104.8 G 15.5 T $\sim$ 39.7 E
+Latency $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
+Memory $164.2$ GB - -
+TABLE XIII: Computational complexity and speed analysis on the LasHeR dataset.
 Method Base Model Param (M) FLOPs (G) FPS SR PR
@@
 Ours SUTrack 141.1 243.7 24.8 62.9 77.8
-Figure 10: Visual comparisons of the tracking performance of different methods on the (Left) RGB-Event, (Middle) RGB-Depth and (Right) RGB-Thermal datasets.
 IV-D Computational Efficiency
 In addition to tracking accuracy, we evaluate the computational efficiency of our sensitivity-aware regularization tuning by analyzing both training and inference overhead.
-Efficiency of Regularized Fine-tuning.
+Computational Overhead of Prior Sensitivity.
+Prior sensitivity is estimated via a group-wise FIM approximation with eigen-decomposition.
+With the operation-diagonal (group-wise) approximation and Rayleigh-quotient probing, we reduce the computational cost from intractable to a practical level.
+We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table XIV to quantify its efficiency.
+While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior sensitivity estimation introduces additional, non-trivial computation.
+Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference.
+Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle (see Appendix for details).
+Efficiency of Sensitivity-Regularized Fine-tuning.
 We compare the training efficiency of our method against vanilla fine-tuning.
@@
 Inference Cost and Speed.
-We compare the computational complexity and runtime efficiency of existing methods, focusing on models built upon OSTrack, DropTrack and SUTrack.
-Note that the proposed regularization techniques are applied only during training, imposing no additional computational burden during testing.
-All experiments are conducted on the same system, which includes an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and one NVIDIA 3090Ti GPU.
-As reported in Tab.
-XII, our method supports real-time tracking at 24.8 frames per second, while delivering top-tier performance.
-When integrated into ViPT, our approach (“ViPT+Ours”) yields both superior speed and performance.
-Under the DropTrack base model, it still operates at 48.8 frames per second, maintaining a respectable speed given its focus on accuracy.
-IV-E Visual Comparison
-Some representative visualization results are illustrated in Fig.
-10.
-For clarity, all visual instances are explicitly labeled by the predicted bounding box.
-Horizontally, each instance is represented by a pair of pictures - the right image shows the main view, while the left displays the auxiliary modality.
-Our method demonstrates a marked improvement over state-of-the-art multi-modal methods, especially in handling complex scenarios where cross-modal integration plays a crucial role.
-It excels in tracking in conditions like motion blur, occlusions, and low illumination, where other methods struggle.
-These results demonstrate that our method improves tracking accuracy and provides a more stable, efficient way to integrate auxiliary data.
+We compare the computational complexity and runtime efficiency of representative baselines, focusing on trackers instantiated with three widely used pre-trained backbones: OSTrack, DropTrack and SUTrack.
+For a fair comparison, we report both theoretical cost (e.g., parameter count and FLOPs) and practical throughput (FPS) under a unified evaluation protocol.
+All experiments are conducted on the same hardware platform, equipped with an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and a single NVIDIA RTX 3090Ti GPU.
+As reported in Table XIII, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting.
+When integrated into ViPT (“ViPT+Ours”), it yields both superior speed and performance.
+Under the DropTrack base model, the resulting tracker still operates at 48.8 FPS, maintaining a respectable speed given its focus on accuracy.
+IV-E Limitations and Discussion
+From a computational-efficiency perspective, a practical consideration is that prior sensitivity estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles.
+Additionally, the formulation defines the Fisher expectation over the pre-training dataset, so applying the method as-is may require access to (or a governed surrogate for) pre-training data—an assumption that does not always hold under licensing, privacy, or retention constraints—potentially limiting portability across model sources and deployment settings.
+Addressing these constraints, e.g., by exploring data-free or proxy-based sensitivity estimation—remains an important direction for future work.
 V Conclusion
@@
 We believe these insights will pave the way for further advancements in multi-modal transfer learning, particularly in the context of scene perception tasks.
-Proof of Proposition 1.
-  1. 1.
+Appendix A Proof of Derivations
+A-A Proof of Proposition 1
+  1.
 Bounded FIM Error.
@@
     In particular, under the common assumption that the top-$K$ eigenvalues in each group capture the most significant curvature, where $\mathrm{Tr}(\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})})\ll\mathrm{Tr}(\mathcal{F}^{(\theta_{0})})$, the residual eigenvalue mass is small.
-  2. 2.
+  2.
 Bounded Generalization Gap Error.
@@
     In practice, by retaining the top-$K$ eigenvalues in each FIM block (which capture the majority of the FIM energy) and replacing the remaining eigenvalues with an isotropic average, one obtains a low-rank Fisher matrix approximation that preserves the important Riemannian geometry of the parameter space while bounding the distortion in any distance or generalization metric by the residual eigenvalue mass.
+A-B Proof of Transfer Sensitivity
+We state the local assumptions under which the transfer sensitivity bound in Eq. (9)-(12) holds and provide a more explicit derivation.
+Assumptions.
+Lipschitz-continuous gradients.
+The loss $L(\theta\mid M)$ is differentiable and $\beta$-smooth in a neighborhood of the current $\theta$, i.e., $\|\nabla L(\theta+\Delta\mid M)-\nabla L(\theta\mid M)\|_{2}\leq\beta\|\Delta\|_{2}$.
+This implies a first-order Taylor expansion with a quadratic remainder: $L(\theta+\delta\mid M)=L(\theta\mid M)+G^{\top}\delta+r(\delta)$ with $|r(\delta)|\leq\frac{\beta}{2}\|\delta\|_{2}^{2}$.
+Bounded single-step gradient magnitude.
+The per-step gradient magnitude is bounded (a standard assumption in stochastic optimization), which is used when relating gradient sparsity to the amplification of $\|G\|_{2}$ under a fixed $\|G\|_{1}$.
+Explicit Perturbation model. $\delta$ and $\delta^{\prime}$ are i.i.d. perturbations around the current update step (as in Eq. (10)).
+Setup.
+Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall
+$\epsilon_{\text{ada}}=\mathbb{E}\left[|S\left(\theta,\delta\mid M\right)-S\left(\theta,\delta^{\prime}\mid M\right)\right|],$ (17)
+$S(\theta,\delta\mid M)=\mathcal{L}(\theta\mid M)-\mathcal{L}(\theta+\delta\mid M).$ (18)
+We model perturbations around a single update step as
+$\delta=\alpha(\mathcal{\mathcal{G}}+\xi),\qquad\delta^{\prime}=\alpha(\mathcal{G}+\xi^{\prime}),\qquad$ (19)
+where $\xi,\xi^{\prime}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}I)$, so that $\delta$ and $\delta^{\prime}$ share the same mean step $\alpha\mathcal{G}$ but differ by stochastic perturbations.
+Linearization ( Lipschitz continuity).
+Assume $L(\theta\mid M)$ is differentiable and locally $\beta$-smooth around $\theta$, i.e., $\|\nabla\mathcal{L}(\theta+\delta\mid M)-\nabla\mathcal{L}(\theta\mid M)\|_{2}\leq\beta\|\delta\|_{2}$.
+Then a first-order Taylor expansion yields
+$\mathcal{L}(\theta+\delta\mid M)=\mathcal{L}(\theta\mid M)+\mathcal{G}^{\top}\delta+r(\delta),\quad|r(\delta)|\leq\frac{\beta}{2}\|\delta\|_{2}^{2},$ (20)
+and hence
+$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta).$ (21)
+Therefore,
+$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})).$ (22)
+When perturbations are sufficiently small (so the remainder term is dominated), the leading term is
+$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi).$ (23)
+A guaranteed upper bound.
+Let $X:=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi)$.
+Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus
+$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (24)
+Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized sensitivity:
+$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (25)
+This shows that $\epsilon_{\text{ada}}$ scales at most linearly with $\|\mathcal{G}\|_{2}$.
+Tighter closed form.
+Moreover, for a zero-mean Gaussian $X\sim\mathcal{N}(0,s^{2})$, $\mathbb{E}|X|=s\sqrt{2/\pi}$. Hence,
+$\mathbb{E}|X|=\sqrt{\frac{2}{\pi}}\sqrt{2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}}=\frac{2\alpha\sigma}{\sqrt{\pi}}\,\|\mathcal{G}\|_{2}\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (26)
+where the last inequality holds since $2/\sqrt{\pi}\leq\sqrt{2}$.
+Deviation from linearization.
+From Eq. (22)–(20), the approximation error is controlled by the Taylor remainders:
+$\Big|\epsilon_{\text{ada}}-\mathbb{E}|X|\Big|\leq\mathbb{E}|r(\delta)-r(\delta^{\prime})|\leq\frac{\beta}{2}\mathbb{E}\big(\|\delta\|_{2}^{2}+\|\delta^{\prime}\|_{2}^{2}\big)=\beta\,\mathbb{E}\|\delta\|_{2}^{2},$ (27)
+so the linear scaling is tightest in locally smooth regions and for sufficiently small perturbations.
+Appendix B Additional ablation studies
+B-A Computational Overhead of Prior Sensitivity under Larger Backbones
+In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior sensitivity estimation under larger backbones, such as ViT-L and Swin-L.
+Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Tab.
+XIV.
+It is important to note that the total cost of the offline FIM process is data-dependent, scaling with the size of the pre-training dataset and the number of iterations required for convergence given the model capacity.
+Consequently, rather than providing a variable total estimate, we report the deterministic per-group and per-iteration costs to accurately reflect the computational costs.
+TABLE XIV: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
+Model Metric Granularity Computational Cost
+Per-Group (Operation-wise) Per-Iteration (Model-wise)
+ViT-L FLOPs Theoretical Ops 170.2 GFLOPs 49.4 TFLOPs
+Latency Wall-clock Time 452.5 ms 131.2 s
+[…]
+Swin-L FLOPs Theoretical Ops 323.1 GFLOPs 93.7 TFLOPs
+Latency Wall-clock Time 687.8 ms 206.3 s
+[…]
+B-B Training Efficiency of SRFT
+We explicitly acknowledge that unlike plug-and-play PEFT methods (e.g., Prompt Tuning, Adapters, LoRA), which incur zero preprocessing cost via random initialization, our SRFT involves an offline FIM estimation phase.
+However, we argue that this offline computation represents a strategic “pay once, benefit long-term” investment.
+This upfront cost is justified by the following characteristics over PEFT:
+1.
+Accelerated Convergence: The FIM estimation is a one-shot process performed on the pre-training datasets.
+The resulting sensitivity map is then frozen and reused for all subsequent multi-modal tracking tasks.
+Furthermore, the pre-trained weight acts as a “smart initialization” compared to the random initialization of PEFT.
+By guiding the optimization along the most reliable directions from the start, SRFT significantly accelerates convergence.
+As shown in Tab.
+XV and Fig.
+11, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
+2.
+Zero Inference Latency: This is of paramount importance for real-time applications.
+As detailed in the original manuscript, the proposed regularization techniques are applied exclusively during training.
+Consequently, they impose no storage overhead and zero additional inference latency during testing, which is crucial for real-time tracking scenarios.
+3.
+Superior Performance: Finally, the computational cost translates directly into performance gains.
+By leveraging the global sensitivity information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
+TABLE XV: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods.
+All experiments are conducted based on the pre-trained OSTrack-B256.
+“Preprocessing” refers to prior sensitivity estimation.
+The training epochs (listed sequentially from left to right) correspond to the FE108, VisEvent, CoeSot, LasHeR, and DepthTrack datasets, respectively.
+Note that UnTrack employs a joint training strategy on the VisEvent, LasHeR, and DepthTrack datasets.
+Method Trainable Parameter (M) Preprocess Time (h) Total Epochs Training Time (h) Total Time (h)
+ViPT (Prompt) 0.8 - 300 $(60\times 5)$ $15.3min\times 300\approx 76.5\text{ h}$ 76.5
+ViPT + SRFT 86.4 47.8 h 100 $(20\times 5)$ $15.9min\times 100\approx 26.5\text{ h}$ 74.3
+SDSTrack (Adapter) 14.8 - 205 $(50\times 3+40+15)$ $21.9min\times 205\approx 74.8\text{ h}$ 74.8
+SDSTrack + SRFT 100.6 47.8 h 85 $(20\times 3+15+10)$ $22.8min\times 85\approx 32.3\text{ h}$ 80.1
+UnTrack (LoRA) 6.6 - 80 $18.5min\times 80\approx 19.3\text{h}$ 19.3
+SDSTrack + SRFT 92.4 47.8 h $40$ $19.4min\times 40\approx 10.3\text{ h}$ 58.1
+Figure 11: Comparison of training loss for different multi-modal tracking methods on DepthTrack dataset.
+Note that all methods share the same loss function and adopt pre-trained OSTrack initialization.
+Evidently, our method converges in fewer training iterations and reaches a lower loss value, effectively mitigating underfitting.
+Moreover, the training dynamics are noticeably more stable, exhibiting reduced oscillation and improved consistency throughout fine-tuning.
+Figure 12: Schematic illustration of different sensitivity fusion scheduling strategies.
+We compare the evolution of the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
+B-C Different Sensitivity Fusion Scheduling Strategies
+Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain).
+In the paper, we explicitly define a dynamic linear schedule to harmonize the two sensitivities: at the beginning, the prior sensitivity has weight $\kappa$, and as training proceeds the transfer sensitivity gradually increases until they reach the same weight $\kappa$ at the end.
+This is formalized by Eq.
+(14), where the combined sensitivity is a convex combination with linear dependence on training progress $\frac{t}{T}$.
+As summarized in the Fig.
+12 and Tab.
+XVI, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under sensitivity harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$.
+To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training.
+Overall, the linear schedule is consistently best or tied-best.
+When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime.
+However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile.
+Mechanistically, exponential/cosine may over-emphasize one sensitivity at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization.
+In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
+TABLE XVI: Comparison of different sensitivity fusion schedules on FE108 dataset.
+We compare the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
+Schedule Prior Sensitivity Weight Transfer Sensitivity Weight $\kappa$ SR PR
+Exponential $w^{p}=\kappa\left(\frac{1-\kappa}{\kappa}\right)^{\frac{t}{T}}$ $w^{t}=1-w^{p}$ 0.3 65.7 94.2
+0.6 67.2 96.3
+0.9 66.4 94.1
+Cosine $w^{p}=(1-\kappa)+\frac{2\kappa-1}{2}\left(1+\cos\left(\frac{t\pi}{T}\right)\right)$ $w^{t}=1-w^{p}$ 0.3 66.6 95.0
+0.6 67.3 96.4
+0.9 66.7 95.2
+Piecewise (4-Stage) $w^{p}=\kappa+\frac{1-2\kappa}{3}\lfloor\frac{4t}{T}\rfloor$ $w^{t}=1-w^{p}$ 0.3 66.8 94.9
+0.6 67.1 96.0
+0.9 66.5 95.3
+Linear (Ours) $w^{p}=\kappa+(1-2\kappa)\frac{t}{T}$ $w^{t}=1-w^{p}$ 0.3 66.9 95.1
+0.6 67.4 96.5
+0.9 66.9 95.5
+Figure 13: Visual comparisons of the tracking performance of different methods on the (Left) RGB-Event, (Middle) RGB-Depth and (Right) RGB-Thermal datasets.
+B-D Visual Comparison
+Some representative visualization results are illustrated in Fig.
+13.
+For clarity, all visual instances are explicitly labeled by the predicted bounding box.
+Horizontally, each instance is represented by a pair of pictures - the right image shows the main view, while the left displays the auxiliary modality.
+Our method demonstrates a marked improvement over state-of-the-art multi-modal methods, especially in handling complex scenarios where cross-modal integration plays a crucial role.
+It excels in tracking in conditions like motion blur, occlusions, and low illumination, where other methods struggle.
+These results demonstrate that our method improves tracking accuracy and provides a more stable, efficient way to integrate auxiliary data.

### v3

@@
-Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning
+
 Abstract
 This paper tackles the critical challenge of optimizing multi-modality trackers by effectively adapting pre-trained models for RGB data.
-Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off.
-To mitigate this dilemma, we propose a novel sensitivity-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities.
-Through a comprehensive investigation of the transition from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are the primary drivers of this issue.
-Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization.
-Subsequently, we characterize transfer sensitivities during the tuning phase, emphasizing adaptability and stability.
-By incorporating these sensitivities as unified regularization terms, our method significantly enhances the transferability across modalities.
-Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modality tracking benchmarks.
-The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
+Existing fine-tuning paradigms oscillate between excessive flexibility and over-restriction, both leading to suboptimal plasticity-stability trade-offs.
+To mitigate this dilemma, we propose a novel significance-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter significance.
+Through a comprehensive investigation of the transition from pre-trained to multi-modality contexts, we identify that parameters crucial to preserving foundational patterns and managing cross-domain shifts are the primary drivers of this issue.
+Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior significance, dedicated to preserving generalization.
+Subsequently, we characterize transfer significance during the fine-tuning phase, emphasizing adaptability and stability.
+By incorporating these parameter significance terms as unified regularization, our method markedly enhances transferability across modalities.
+Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modal tracking benchmarks.
+The source code and models are publicly available at https://github.com/zhiwen-xdu/SRTrack.
+journal: International Journal of Computer Vision
 Figure 1: Optimization trajectory analysis on LasHeR.
@@
 As visualized by the colorful arrows, our method effectively mitigates the misfitting issue and enhances multi-modality trackers with superior generalization and stability.
-I Introduction
-Object tracking, a foundation task of visual perception, has seen significant advancements over the past decades [16, 54, 67, 4].
+1 Introduction
+Object tracking, a foundation task of visual perception, has witnessed remarkable advancements over the past decades (Hong et al., 2024; Xie et al., 2024; Zheng et al., 2024a; Cai et al., 2024).
 Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions.
-Therefore, multi-modality tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest.
-With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [58, 29, 10, 9].
-There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
-Some existing approaches have explored the full fine-tuning (FFT) paradigm [45, 73, 41], where trackers are initialized with pre-trained weights and subsequently optimized for task-specific objectives.
-Nevertheless, due to the significant domain gaps and limited scale of auxiliary modalities, these methods struggle to preserve the pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance).
-In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [18, 21, 7].
+Therefore, multi-modality tracking with more comprehensive sensory signals (e.g., event, depth, thermal) has garnered growing interest.
+With the popularity of the data-driven methods in the object tracking community, both data scale and model capacity have experienced huge explosions in recent years (Ye et al., 2022; Lin et al., 2022; Chen et al., 2023, 2025b).
+There is a prevailing paradigm that explores these pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence by starting from well-structured pre-trained weights.
+Concretely, some existing approaches have explored the full fine-tuning (FFT) paradigm (Wang et al., 2023; Zhu et al., 2023c; Sun et al., 2025), where multi-modality trackers are initialized with RGB-based weights and subsequently optimized for task-specific objectives.
+Nevertheless, while FFT offers maximum flexibility and swiftly adapts to the target domain, substantial domain gaps coupled with limited scale of auxiliary modalities, hinder the retention of pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance).
+In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) (Hu et al., 2021; Jia et al., 2022; Chen et al., 2022).
 The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modality-specific ones to retain prior knowledge.
-Several methods fall under this umbrella [70, 17, 51], including prompt tuning, visual adapter, etc.
+Several methods fall under this umbrella (Zhu et al., 2023a; Hou et al., 2024; Wu et al., 2024b), including prompt tuning, visual adapter, etc.
 Although effective, PEFT-based methods impose rigid constraints on the primary model weights, resulting in underfitting (i.e., a restricted upper bound on training performance) when handling the significant distribution drifts.
-In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as Fig.
+In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting (i.e., overfitting and underfitting) and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as clearly illustrated in Fig.
 1).
 In this work, we endeavor to mitigate the misfitting dilemma in cross-modal tracker adaptation by strategically regularizing the learning process.
-By probing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability.
-To mitigate this, we propose a sensitivity-regularized fine-tuning framework that delicately calibrates the gradient updates to boost model transfer with precision.
+By analyzing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter significance, in its various aspects, is the primary cause of degraded prior generalizability and transfer adaptability.
+To mitigate this, we propose a significance-regularized fine-tuning (SRFT) framework that delicately calibrates the gradient updates to boost model transfer with precision.
 Specifically, we optimize multi-modality trackers from the following perspectives.
-  1. (1)
-    Formulating Prior Sensitivity.
+  1. 1.
+    Formulating Prior Significance.
 We commence by investigating the tangent space of pre-trained parameters, utilizing it as a critical indicator to assess and preserve prior generalizability for downstream fine-tuning.
-Furthermore, we estimate this sensitivity via an eigen-decomposition approximation.
-  2. (2)
-    Modeling Transfer Sensitivity.
-To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate tuning instability.
-Accordingly, we formulate transfer sensitivity using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
-  3. (3)
-    Sensitivity Regularized Tuning.
-By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, sensitivity-driven updates.
-This mechanism facilitates seamless transfer across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
-Our method strategically guides the cross-domain fine-tuning process to optimize downstream tracking tasks.
+Furthermore, we estimate this significance via an eigen-decomposition approximation.
+  2. 2.
+    Modeling Transfer Significance.
+To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate fine-tuning instability.
+Accordingly, we formulate transfer significance using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
+  3. 3.
+    Significance Regularized Tuning.
+By harnessing the identified parameter significance, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, significance-driven updates.
+This mechanism facilitates seamless fine-tuning across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
+Our method strategically guides the cross-domain fine-tuning process to optimize downstream multi-modality tracking tasks.
 Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modality tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models.
-Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
+Comprehensive ablation studies and parameter significance measurements confirm the effectiveness of the significance-aware regularization fine-tuning strategy.
 In summary, the main contributions of this paper are:
-  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
-  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
+  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework (SRFT) to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
+  • we formulate the parameter significance with respect to pre-trained knowledge and transfer stability, and introduce a significance-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
   • we conduct comprehensive experiments covering three multi-modality tracking tasks and seven benchmarks using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels.
 The remainder of the paper is organized as follows.
-Section II reviews existing literature on multi-modal trackers and cross-modal transfer learning methods.
-In Section III, we rigorously formulate prior and transfer sensitivities from both pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-regularized tuning.
-Section IV presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs.
-Finally, Section V concludes the paper.
-II Related Work
-II-A Multi-Modal Object Tracking
-Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [48, 2, 28, 68].
-However, the unreliability of RGB-only paradigms to adverse conditions has catalyzed research into multi-modal tracking, where auxiliary signals (e.g., event, depth, thermal data) are integrated to remedy the intrinsic deficiencies of visible imagery [66].
-For instance, event cameras provide robust dynamic scene information under extreme motion or lighting.
-Leveraging this, [72, 11, 46, 63, 43, 44] combined RGB and event streams to predict objects in high-speed and low-dynamic scenarios.
-Meanwhile, depth cues offer geometric priors that are particularly useful for handling occlusion and background clutter.
-[32, 36, 56, 31] exploited this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes.
-Similarly, thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination.
-[62, 19, 27, 52] demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
-In summary, contemporary methods emphasized effective multi-modal interaction and fusion [42, 45, 59, 65].
-With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [58, 49, 9] have demonstrated remarkable generalization across diverse scenarios.
-These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers.
+Section 2 reviews existing literature on multi-modal trackers and cross-modal transfer learning methods.
+In Section 3, we rigorously formulate prior and transfer parameter significance from both pre-training and fine-tuning perspectives, followed by the implementation of significance-regularized tuning.
+Section 4 presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs.
+Finally, Section 5 concludes the paper.
+2 Related Work
+2.1 Multi-Modal Object Tracking
+Object tracking involves localizing an object across frames given its initial appearance (Wei et al., 2023; Bai et al., 2024; Lin et al., 2024; Zheng et al., 2024b).
+However, the vulnerability of RGB-only trackers under adverse conditions has driven the development of multi-modal tracking, where auxiliary cues complement the intrinsic deficiencies of visible imagery (Zhang et al., 2023c).
+For instance, event cameras provide robust dynamic information under extreme motion or lighting, enabling RGB-event fusion for high-speed and low-dynamic tracking (Zhu et al., 2022; Chen et al., 2024b; Wang et al., 2024; Zhang et al., 2024b; Wang et al., 2025a, b).
+Depth offers geometric priors that help handle occlusion and background clutter, and has been integrated to improve robustness in crowded scenes (Lukezic et al., 2019; Qian et al., 2021; Yan et al., 2021b; Liu et al., 2025).
+Thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination.
+(Zhang et al., 2021b; Hui et al., 2023; Li et al., 2025; Xiang et al., 2025) demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
+In summary, contemporary methods emphasized effective multi-modal interaction and fusion (Zhang et al., 2024a, b; Chen et al., 2024a; Tan et al., 2025a, b; Feng et al., 2025).
+With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers (Ye et al., 2022; Wu et al., 2023; Chen et al., 2025b) have demonstrated remarkable generalization across diverse scenarios.
+These advances have shifted the paradigm toward transferring pre-trained models for designing high-performance multi-modal trackers.
 Consequently, multi-modal tracking is increasingly driven by reusing RGB-induced semantic priors, particularly when annotated multi-modal data are scarce or noisy.
-In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modality transfer.
-II-B Cross-modal Transfer Learning
+In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modal transfer learning.
+2.2 Cross-modal Transfer Learning
 To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged.
-Some works adhere to the full fine-tuning (FFT) paradigm [42, 45, 73, 41], where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks.
+Some works adhere to the full fine-tuning (FFT) paradigm (Tang et al., 2022; Wang et al., 2023; Zhu et al., 2023c; Sun et al., 2025), where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks.
 These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model.
-Representatively, SUTrack [9] employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization.
+Representatively, SUTrack (Chen et al., 2025b) employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization.
 While effective, one primary dilemma may be innate to FFT: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation.
-This imbalance frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data.
-To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [21, 7, 18], the field has pivoted toward parameter-efficient fine-tuning (PEFT).
+This mismatch frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data.
+To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities (Jia et al., 2022; Chen et al., 2022; Hu et al., 2021), the field has pivoted toward parameter-efficient fine-tuning (PEFT).
 The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task.
 By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting.
-For example, ProTrack [57] and ViPT [70] modulated RGB features by introducing trainable auxiliary tokens into attention layers.
-Similarly, methods like BAT [5] and SDSTrack [17] inserted lightweight adapter modules between attention layers for cross-modal shift compensation.
-More recently, unified trackers like UnTrack [51] and OneTracker [16] employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs.
+For example, ProTrack (Yang et al., 2022) and ViPT (Zhu et al., 2023a) modulated RGB features by introducing trainable auxiliary tokens into attention layers.
+Similarly, methods like BAT (Cao et al., 2024) and SDSTrack (Hou et al., 2024) inserted lightweight adapter modules between attention layers for cross-modal shift compensation.
+More recently, unified trackers like UnTrack (Wu et al., 2024b) and OneTracker (Hong et al., 2024) employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs.
 Despite their effectiveness, PEFT methods suffer from intrinsic limitations.
@@
 Remark.
-While recent works have explored sensitivity-aware mechanisms in efficient fine-tuning (SPT) and continual learning (CL), our SRFT introduces a fundamentally distinct paradigm in objectives and mechanisms.
-Specifically, methods like SPT [15] and CDRA-SPT [8] utilized sensitivity to prioritize structural adaptation, such as parameter masking or dynamic rank allocation for parameter efficiency.
-Conversely, MACL [47] employed sensitivity to prevent forgetting in sequential tasks by minimizing loss variance.
-Crucially, these approaches typically operate with an exclusive focus on the target or current data.
-Diverging from this, SRFT establishes a gradient-regularized paradigm.
-We formulate sensitivity to balance the preservation of pre-trained knowledge with sparse transfer gradient, via a hybrid sensitivity from pre-trained to multi-modal contexts.
-This distinction renders our method orthogonal to existing FFT, PEFT, and SPT frameworks.
-Figure 2: Network architecture of the multi-modality trackers.
+Beyond fine-tuning strategies in multi-modal tracking, we also review related fine-tuning paradigms from adjacent domains that could be conflated with ours, to more clearly distinguish our contribution.
+In particular, some recent studies have explored parameter sensitivity in efficient fine-tuning (SPT) and continual learning (CL).
+Notably, our method differs fundamentally from these approaches in both objective and mechanism.
+Specifically, SPT (He et al., 2023) and CDRA-SPT (Chen et al., 2025a) leveraged parameter sensitivity to guide structural adaptation for efficiency, e.g., sparse parameter updates or dynamic rank allocation.
+MACL (Wang and Huang, 2024) used sensitivity to mitigate forgetting in sequential tasks by minimizing loss variance.
+Crucially, these approaches typically operate with an exclusive focus on the target context.
+Diverging from this, SRFT adopts a gradient-regularized paradigm: we define a hybrid parameter significance spanning pre-trained and cross-modal domains to balance prior generalization with target adaptability.
+This distinction renders our method orthogonal to existing FFT, PEFT, and SPT methods.
+Figure 2: Network architecture of our multi-modality trackers.
 All modules are initialized with the weights of a pre-trained RGB tracker.
-We only fine-tune backbones and fusion blocks with sensitivity-aware regularization terms.
-III Proposed Method
+We only fine-tune backbones and fusion blocks with significance-aware regularization terms.
+3 Proposed Method
 Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers.
 To unlock the full potential of pre-trained trackers, we revisit the core principles governing the cross-modal transfer process.
-In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning.
-Next, Section III-B identifies and quantifies pronounced prior and transfer sensitivities that characterize pre-trained knowledge and cross-domain adaptation.
-Finally, in Section III-C, we incorporate these sensitivities into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
-III-A Preliminaries
+In Section 3.1, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning.
+Next, Section 3.2 identifies and quantifies pronounced prior and transfer parameter significance that characterize pre-trained knowledge and cross-domain adaptation.
+Finally, in Section 3.3, we incorporate these significance measurements into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
+3.1 Preliminaries
 Network Architecture of Multi-modality Trackers.
 Fig.
-2 depicts the overall architecture of our multi-modal tracker, in which all modules are initialized with the weights of a pre-trained RGB tracker.
-For a multi-modal tracker, the RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens.
+2 depicts the architecture of our multi-modal trackers, in which all modules are initialized with the weights of a pre-trained RGB tracker.
+For the multi-modal tracker, RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding template and search tokens.
 Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction.
-Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens.
+Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks (e.g., layers 2, 5, 8, and 11) for multi-stage fusion by concatenating multi-modal template and search tokens.
 Finally, the fused features are fed into the box head to estimate the object state.
 To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
+Additional model details can be found in Appendix A.
 Learning Process of Cross-Domain Transfer.
 We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameters $\theta$.
-It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box.
-In this paper, we focus on efficiently transferring the models pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$.
-In this setting, let $\theta_{0}$ be the pre-trained parameters (optimal on $D_{0}$), which serve as initialization.
-After fine-tuning on $D_{t}$, the parameters become $\theta_{t}$.
-Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
+The model takes a task-specific dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x_{i}$ represents a multi-modal input pair and $y_{i}$ is the corresponding object bounding box or output for the task.
+In this work, we focus on efficiently transferring the models that have been pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded versions of images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$.
+Let $\theta_{0}$ represent the pre-trained parameters, which are optimal for the source domain $D_{0}$, and these parameters are used as the weight initialization.
+After fine-tuning the model on the target domain $D_{t}$, the model parameters become $\theta_{t}$.
+Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is formulated as follows:
 $\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\quad\theta^{(0)}=\theta_{0},$ (1)
@@
 1.
-This motivates a more principled transfer learning approach, as described next.
+This occurs because the model tends to over-adapt to the unstable target domain $D_{t}$, while neglecting the generalization capabilities learned from the source domain $D_{0}$.
+These challenges motivate a more principled transfer learning approach, as described next.
 Figure 3: Loss-parameter manifold schematic across different fine-tuning paradigms.
@@
 In contrast, our SRFT performs optimization within an approximated pre-trained tangent space, indicating a better plasticity-stability trade-off.
-III-B Modeling Sensitivity for Multi-Modality Tracking
-Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) [34].
-We regulate this trade-off via two intrinsic parameter sensitivities, identifying the evolving parameter importance from pre-trained to target tasks: prior sensitivity, reflecting its importance to pre-trained knowledge, and transfer sensitivity, indicating its role in target-task adaptation.
-Excessive changes to highly sensitive parameters in either aspect can hinder the transfer process.
-Manifold hypothesis.
+3.2 Modeling Significance for Multi-Modality Tracking
+Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) (Mermillod et al., 2013; Zheng et al., 2025; Zhou et al., 2025).
+We regulate this trade-off via two intrinsic parameter significance, identifying the evolving parameter importance from pre-trained to target tasks: prior significance, reflecting its importance to pre-trained knowledge, and transfer significance, indicating its role in target-task adaptation.
+Excessive changes to highly significant/fragile parameters in either aspect can hinder the transfer process.
+Loss-Parameter Manifold Hypothesis.
 As shown in Fig.
-3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$.
-Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ [13, 40].
-Based on the aforementioned manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig.
+3, the parameter model set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$.
+Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ Bengio et al.
+(2013); Fefferman et al.
+(2016); Song and Ermon (2019); Meilă and Zhang (2024).
+Based on the manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig.
 3.
 Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
-III-B1 Safeguarding Generalization via Prior Sensitivity
-In this context, we first analyze the influence of the tangent space on model generalization from the pre-trained loss-parameter manifold, and then design an efficient eigen-decomposition method to quantify this sensitivity.
-Exploring Parameter Manifold Geometry. We first define the joint empirical risks for tuning models over the union of pre-training and downstream tasks as
+3.2.1 Safeguarding Generalization via Prior Significance
+In this context, we first analyze the influence of the tangent space on model generalization from the pre-trained loss-parameter manifold, and then design an efficient eigen-decomposition method to quantify this prior significance.
+Exploring Loss-Parameter Manifold Geometry. We first define the joint empirical risks for tuning models over the union of pre-training and downstream tasks as:
 $\mathcal{L}\left(\theta\mid D_{u}\right)=\mathcal{L}\left(\theta\mid D_{t}\right)+\beta\mathcal{L}\left(\theta\mid D_{0}\right),$ (2)
-where $D_{u}=D_{0}\bigcup D_{t}$, $\beta>0$ balances the contributions of pre-trained task, and $\mathcal{L}\left(\theta\mid D_{0}\right)$ serves as a regularization term to refine the vanilla objective.
+where $D_{u}=D_{0}\bigcup D_{t}$, $\beta>0$ balances the contributions of pre-trained task, and $\mathcal{L}\left(\theta\mid D_{0}\right)$ serves as a regularization term to refine the vanilla objective $\mathcal{L}\left(\theta\mid D_{t}\right)$.
 Under this joint objective, the pre-trained weights $\theta_{0}$ minimize $\mathcal{L}\left(\theta\mid D_{0}\right)$ and can serve as a local optimum of $\mathcal{L}\left(\theta\mid D_{u}\right)$.
@@
 $\frac{1}{2}\left(\theta-\theta_{0}\right)^{T}\mathcal{F}^{\left(\theta_{0}\right)}\left(\theta-\theta_{0}\right)+\mathcal{O}\left(\left\|\theta-\theta_{0}\right\|^{3}\right),$
-where $\nabla\mathcal{L}\left(\theta_{0}\right)\approx\mathbf{0}$, and $\mathcal{F}^{\left(\theta_{0}\right)}\in\mathbb{R}^{|\theta|\times|\theta|}$ is expectation of Hessian matrix $\frac{\partial^{2}\mathcal{L}}{\partial\theta_{i}\theta_{j}}$ over dataset $D_{0}$, also called as the Fisher Information Matrix (FIM) [1] at $\theta_{0}$.
+where $\nabla\mathcal{L}\left(\theta_{0}\right)\approx\mathbf{0}$, and $\mathcal{F}^{\left(\theta_{0}\right)}\in\mathbb{R}^{|\theta|\times|\theta|}$ is expectation of Hessian matrix $\frac{\partial^{2}\mathcal{L}}{\partial\theta_{i}\theta_{j}}$ over dataset $D_{0}$, also called as the Fisher Information Matrix (FIM) Amari et al.
+(2019) at $\theta_{0}$.
 Principally, it captures the pre-trained tangent space, along which the impact on the pre-trained loss is minimized.
@@
 This generalization gap is measured as the weight distance within the Riemannian manifold defined by FIM.
-Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [30, 50].
-Therefore, FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
-Figure 4: Operation-wise prior sensitivity (i.e., eigen-decomposition FIM map) of the pre-trained OSTrack.
-The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
+Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation (Liu et al., 2022; Wu et al., 2024a).
+Therefore, FIM serves as a natural and principled prior significance to reflect the erosion of pre-trained knowledge.
+Figure 4: Operation-wise prior parameter significance (i.e., eigen-decomposition of the FIM) for the pre-trained OSTrack on the source datasets.
+The high prior significance indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
 Eigen-decomposition Approximation of FIM.
@@
 Parameters from different groups thus span orthogonal subspaces with no second-order coupling.
-Empirical studies [14, 37] have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs.
+Empirical studies (Ghorbani et al., 2019; Rame et al., 2022) have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs.
 Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
-We obtain these leading eigenpairs via Rayleigh-quotient probing [26]. For any non-zero direction $\epsilon$, the Rayleigh quotient is defined as:
+We obtain these leading eigenpairs via Rayleigh-quotient probing (Li, 2015). For any non-zero direction $\epsilon$, the Rayleigh quotient is defined as:
 $\mathcal{R}(\mathcal{F}^{\left(\theta_{0}\right)},\epsilon)=\frac{\epsilon^{T}\mathcal{F}^{\left(\theta_{0}\right)}\epsilon}{\epsilon^{T}\epsilon},$ (5)
@@
 This formula links the sharpest directions of generalization degradation to the dominant FIM eigenvalues, providing both theoretical motivation and practical guidance.
-Proposition 1 (Eigen-based Approximation Error Bound)
-Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition of group $j$, with top-$K$ eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ and eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$.
+Proposition 1(Eigen-based Approximation Error Bound.)
+Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition of parameter group $j$, with top-$K$ eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ and eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$.
 We construct the following approximation:
@@
 Then the discrepancy between these distances is bounded: $|\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})-\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})|\leq\frac{1}{2}\|\Delta\theta\|^{2}(\underset{1\leq j\leq N}{\operatorname{\max}}\lambda_{1}^{j})$.
-Proof 1
-See Appendix I-A.
-Prior Sensitivity Measure.
-Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [23].
+Proof
+See Appendix B.
+Prior Significance Measurement.
+Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes (LeVeque, 1998).
 Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to the Rayleigh quotient of FIM.
-This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters.
+This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most significant parameters.
 For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
@@
 This measure provides a principled yet tractable perspective to assess how each group of the model affects generalization.
-By considering the eigenvalue magnitudes from different perturbation directions, we condense the prior sensitivity of the $j^{\text{th}}$ operation into a single scalar by distilling its dominant values:
+By considering the eigenvalue magnitudes from different perturbation directions, we condense the prior significance of the $j^{\text{th}}$ operation into a single scalar by distilling its dominant values:
 $s_{j}^{p}=\frac{1}{K}\sum_{i=1}^{K}\lambda_{i}^{j}.$ (8)
-The prior sensitivity is equivalent to FIM in Eq.
+The prior significance is equivalent to FIM in Eq.
 (6), as presented in Fig.
@@
 We set $K=10$ in our experiments.
-In practice, we observe that the leading eigenvalue alone (e.g., $K=1$) provides a reliable indicator of prior sensitivity.
-Further qualitative results are reported in Section IV-C.
-III-B2 Stabilizing Adaptation via Transfer Sensitivity
+In practice, we observe that the leading eigenvalue alone (e.g., $K=1$) provides a reliable indicator of prior significance.
+Further qualitative results are reported in Section 4.3.
+3.2.2 Stabilizing Adaptation via Transfer Significance
 In this context, we analyze the impact of sparse transfer gradients on model adaptation and propose rebalancing them using the gradient matrix.
 Characterizing Transfer-Gradient Sparsity.
-While prior sensitivity addresses forgetting of pre-trained domain, inadaptability in the downstream domain is notably pronounced.
-Multi-modal tracking often involves diverse and heterogeneous samples exhibiting substantial gaps, such as auxiliary inputs or degraded images.
+While prior significance addresses forgetting of pre-trained domain, inadaptability in the downstream domain remains notably pronounced.
+Specifically, multi-modal tracking often involves diverse and heterogeneous samples exhibiting substantial gaps, such as auxiliary inputs or degraded images.
 These shifts manifest as highly sparse gradients during fine-tuning process on $D_{t}$, as illustrated in Fig.
@@
 1).
-In this context, we investigate the elevated adaptation risk induced by sparse gradients and propose a transfer sensitivity aimed at mitigating its impact.
+In this context, we investigate the elevated adaptation risk induced by sparse gradients and propose a transfer significance aimed at mitigating its impact.
 Formally, we quantify the sparsity of the gradient by comparing its $L_{1}$ and $L_{2}$ norms.
@@
 We define the sparsity measure: $\rho=\frac{\left\|\mathcal{G}\right\|_{1}}{\sqrt{|\theta|}\left\|\mathcal{G}\right\|_{2}}$.
-This definition leverages the fact that $L_{1}$ and $L_{2}$ norms characterize different aspects of the vector: while $L_{1}$ reflects the total variation, $L_{2}$ emphasizes concentration [20].
+This definition leverages the fact that $L_{1}$ and $L_{2}$ norms characterize different aspects of the gradient vector: while $L_{1}$ reflects the total variation, $L_{2}$ emphasizes concentration Hurley and Rickard (2009).
 Thus, their ratio serves as an effective proxy for sparsity structure, where a smaller value means higher sparsity.
-Moreover, noting that single-step gradient magnitudes are typically bounded [35], increasing sparsity (i.e., $\rho:1\rightarrow\frac{1}{\sqrt{|\theta|}}$) can amplify the $L_{2}$ norm:
+Moreover, noting that single-step gradient magnitudes are typically bounded Nesterov (2013), increasing sparsity (i.e., $\rho:1\rightarrow\frac{1}{\sqrt{|\theta|}}$) can amplify the $L_{2}$ norm of the gradient vector:
 $\left\|\mathcal{G}\right\|_{2}=\frac{\left\|\mathcal{G}\right\|_{1}}{\rho\sqrt{|\theta|}}.$ (9)
-Figure 5: Instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning.
-The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
+Figure 5: Instantaneous transfer parameter significance on the VisEvent dataset observed during fine-tuning.
+The high sparsity in the significance map indicates that only a few gradients (light areas) dominate the updates.
 Sparsity-Induced Adaptation Risk.
 We characterize the tuning volatility as the degree to which performance responds to small parameter perturbations during fine-tuning.
-Accordingly, we define the following sensitivity function:
+Accordingly, we define the following response function:
 $S\left(\theta,\delta\mid M\right)=\mathcal{L}\left(\theta\mid M\right)-\mathcal{L}\left(\theta+\delta\mid M\right),$ (10)
@@
 $\varepsilon_{ada}=\underset{\left(\delta,\delta^{\prime}\right)}{\mathbb{E}}\left[|S\left(\theta,\delta\mid M\right)-S\left(\theta,\delta^{\prime}\mid M\right)\right|].$ (11)
-By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq. (11), it can be shown that this risk is proportional to the $L_{2}$ norm of gradient:
+By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq.
+(11), it can be shown that this adaptation risk scales with the $L_{2}$ norm of the gradient vector:
 $\varepsilon_{ada}\leq\alpha\sqrt{2\sigma^{2}}\left\|\mathcal{G}\right\|_{2}.$ (12)
-Derivations details of Eq.
-(9)-(12) can be found in Appendix I-B.
+Derivations details of Eqs.
+(9)-(12) are provided in Appendix B.
 Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk.
 This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
-Transfer Sensitivity Estimation.
-To formalize the estimation of transfer sensitivity, we adapt the formulation of Eq.
-(10) to the standard tuning process.
+Transfer Significance Estimation.
+To formalize the estimation of transfer significance, we adapt the formulation of Eq.
+(10) to the standard fine-tuning process.
 In this case, we adjust the parameter $\theta$ via the gradient $\mathcal{G}$ rather than noise perturbations.
 Therefore, we can derive: $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\mathcal{G}$.
-Here, the sensitivity function is used to assess the overall adjustment of all parameters during transfer.
-To further investigate and regularize gradients, we extend the sensitivity to a parameter-wise formulation:
+Here, the significance function is used to assess the overall adjustment of all parameters during transfer.
+To further investigate and regularize gradients, we extend the transfer significance to a parameter-wise formulation:
 $s_{n}^{t}=\left(\frac{\partial\mathcal{L}\left(\theta\mid M\right)}{\partial\theta_{n}}\right)^{2}.$ (13)
-where $s_{n}^{t}$ denotes the transfer sensitivity of the $n^{\text{th}}$ parameter.
+where $s_{n}^{t}$ denotes the transfer significance of the $n^{\text{th}}$ parameter.
 This formulation allows us to examine the granular impact of parameter changes during the transfer process and rebalance the gradient accordingly.
+Algorithm 1 Significance-Regularized Fine-Tuning
 Input: pre-trained dataset $D_{0}$, downstream task dataset $D_{t}$, pre-trained weight $\theta_{0}$; number of eigenvalues $K$, number of model operations NN, training steps $T$; initialized eigenvalue set $\Lambda$;
 Output: Optimal multi-modal tracker parameters $\theta_{t}$;
-Step1: Prior Sensitivity Estimation:
+Step1: Prior Significance Estimation:
 for batch $(x,y)$ in $D_{0}$ do
@@
 end for
- Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
-Step2: Transfer Sensitivity Estimation and Regularized Tuning:
+ Calculate prior significance $s_{j}^{p}$ via Eq. (8);
+Step2: Transfer Significance Estimation and Regularized Tuning:
 for $i\in\{1,\ldots,T\}$ do
@@
   Compute loss $\mathcal{L}(\theta\mid M_{i})$ and gradients;
-  Update transfer sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
-  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
+  Update transfer significance by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
+  Normalize and combine significance $s_{n}$ by Eq. (14);
   Parameter update via Eq. (15);
 end for
-Algorithm 1 Sensitivity-Regularized Fine-Tuning
-III-C Sensitivity-regularized Tuning
-We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers.
-To align with the unified regularization tuning, we rank both the operation-wise prior sensitivity and parameter-wise transfer sensitivity, and then normalize them within the continuous range of $[0,1]$.
-To harmonize these two sensitivities, we devise a dynamic linear schedule to adjust their weighted combination.
-At the beginning of training, prior sensitivity plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$).
-As training progresses, the influence of transfer sensitivity gradually increases, reaching the same weight $\kappa$ by the end.
+3.3 Significance-Regularized Fine-Tuning
+We regularize the learning process by previously discussed significance metrics to derive enhanced multi-modality trackers.
+To align with the unified regularization tuning, we rank both the operation-wise prior significance and parameter-wise transfer significance, and then normalize them within the continuous range of $[0,1]$.
+To harmonize these two parameter significance, we devise a dynamic linear schedule to adjust their weighted combination.
+At the beginning of training, prior significance plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$).
+As training progresses, the influence of transfer significance gradually increases, reaching the same weight $\kappa$ by the end.
 This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability.
-Formally, at each training step $i$, the combined sensitivity is updated as follows:
+Formally, at each training step $i$, the combined parameter significance is updated as follows:
 $s_{n}=\left(\kappa+(1-2\kappa)\frac{i}{T}\right)s_{j}^{p}+\left(1-\kappa-(1-2\kappa)\frac{i}{T}\right)s_{n}^{t},$ (14)
 where $\theta_{n}\in\theta^{j}$, $T$ is the total number of training steps.
-Finally, we normalize this joint sensitivity within the range of $s_{n}\in[0,0.99]$.
-To this end, during the training process, parameters that are excessively updated will be penalized based on their sensitivity:
-$\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\text{s.t. }\theta^{(0)}=\theta_{0},$ (15)
-where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
-The detailed regularization tuning process is given in Algorithm 1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table VI.
-III-D Learning Objectives
+Finally, we normalize this joint significance within the range of $s_{n}\in[0,0.99]$.
+To this end, during the training process, parameters that are excessively updated will be penalized based on their significance:
+$\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\quad{s.t.}\quad\theta^{(0)}=\theta_{0},$ (15)
+where $\alpha$ is the learning rate.
+This formulation suggests that more significant/sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
+The detailed regularization tuning process is given in Algorithm 1. We set the significance harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table 6.
+Discussion.
+Unlike prior works that rely on static “sensitivity” to rank parameters for sparse tuning, SRFT introduces a fundamentally distinct notion of “significance” — a hybrid, dynamic estimate combining pre-trained Fisher eigen-structure and instantaneous gradient sparsity.
+Instead of selecting the most sensitive parameters for exclusive updating, SRFT prioritizes penalizing them to suppress excessive updates, without structural constraints or discontinuous optimization.
+This regularization-centric design yields a smooth, adaptive update path tailored for cross-modal transfer, diverging from existing sparsity- or retention-oriented approaches.
+Additional comparative experiments can be found in Appendix C.
+3.4 Learning Objectives
 The overall loss function of our method is the same as the foundation model without extra adjustments, shown as:
 $\mathcal{L}=\mathcal{L}_{\text{cls}}+\lambda_{\text{iou }}\mathcal{L}_{\text{iou }}+\lambda_{l_{1}}\mathcal{L}_{1},$ (16)
-where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as [58].
+where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as (Ye et al., 2022).
 Figure 6: Precision and success plots of the FE108, VisEvent, CoeSot and LasHeR datasets. Zoom in to see details.
 Figure 7: Precision scores of different attributes on the VisEvent, CoeSot and LasHeR test sets. Zoom in to see details.
-IV Experiments
-IV-A Experiment Settings
-IV-A1 Benchmark Datasets
+4 Experiments
+4.1 Experiment Settings
+4.1.1 Benchmark Datasets
 To verify the effectiveness and generalization of the proposed method, we conduct comprehensive experiments on seven multi-modal benchmark datasets.
-For RGB-Event tracking, we use FE108 [61] (76 train / 32 test sequences) captured under degraded conditions, VisEvent [45](500 train / 320 test sequences) covering dynamic outdoor scenes, and CoeSot [42] with 578K image-event pairs (824 train / 528 test sequences).
-For RGB-Depth tracking, we adopt DepthTrack [56] (152 train / 50 test videos) and VOT-RGBD2022 [22] (127 test sequences).
-For RGB-Thermal tracking, we use LasHeR [25] (979 train / 245 test sequences) and RGBT234 [24] (234 test videos).
-All training and evaluation strictly follow established protocols [70, 51].
-During fine-tuning, we use the training sets of FE108, VisEvent, and CoeSot for RGB-Event tasks; DepthTrack for RGB-Depth; and LasHeR for RGB-Thermal.
+For RGB-Event tracking, we use FE108 (Zhang et al., 2021a) (76 train / 32 test sequences) captured under degraded conditions, VisEvent (Wang et al., 2023)(500 train / 320 test sequences) covering dynamic outdoor scenes, and CoeSot (Tang et al., 2022) with 578K image-event pairs (824 train / 528 test sequences).
+For RGB-Depth tracking, we adopt DepthTrack (Yan et al., 2021b) (152 train / 50 test videos) and VOT-RGBD2022 (Kristan et al., 2022) (127 test sequences).
+For RGB-Thermal tracking, we use LasHeR (Li et al., 2020) (979 train / 245 test sequences) and RGBT234 (Li et al., 2019) (234 test videos).
+All training and evaluation strictly follow established protocols (Zhu et al., 2023a; Wu et al., 2024b).
+During fine-tuning, we use the training sets of FE108, VisEvent, and CoeSot for RGB-Event tasks; DepthTrack for RGB-Depth tasks; and LasHeR for RGB-Thermal tasks.
 The corresponding test sets, along with VOT-RGBD2022 and RGBT234, are used for evaluation.
-IV-A2 Evaluation Metrics
-Adhering to recognized standards [17, 51, 16, 41], we evaluated tracking performance with the following metrics.
+Table 1: Quantitative comparison on the RGB-Event datasets. The best are marked with “bold”, and the second best are marked with “underline”.
+Method Reference Base Model FE108 VisEvent CoeSot
+SR PR SR PR SR PR
+Image-based Methods (Only-RGB)
+DiMP (Bhat et al., 2019) ICCV’19 ResNet 48.5 71.7 45.2 66.2 56.7 69.1
+Stark-S (Yan et al., 2021a) ICCV’21 ResNet 51.4 75.1 44.6 61.3 55.7 66.7
+OSTrack (Ye et al., 2022) ECCV’22 OSTrack-256 50.7 73.7 53.4 69.6 64.5 77.0
+DropTrack (Wu et al., 2023) CVPR’23 DropTrack-384 52.0 75.9 54.8 72.4 66.1 79.1
+Cross-modal Transfer Learning
+FENet (Zhang et al., 2021a) ICCV’21 DiMP 63.3 92.4 50.8 70.4 53.8 66.1
+AFNet (Zhang et al., 2023a) CVPR’23 DiMP 63.8 92.6 51.8 70.8 52.5 63.8
+CEUTrack (Tang et al., 2022) Arxiv’22 OSTrack-B256 56.7 84.2 55.0 71.6 62.0 76.4
+ProTrack (Yang et al., 2022) ACM MM’22 OSTrack-B256 58.0 85.5 47.1 63.2 66.1 78.4
+HRTrack (Zhu et al., 2023c) ICCV’23 OSTrack-B256 - - - - 63.2 71.9
+ViPT (Zhu et al., 2023a) CVPR’23 OSTrack-B256 65.1 93.5 59.2 75.8 67.7 80.6
+SDSTrack (Hou et al., 2024) CVPR’24 OSTrack-B256 60.0 87.2 59.7 76.7 67.9 80.9
+UnTrack (Wu et al., 2024b) CVPR’24 OSTrack-B256 - - 58.9 75.6 - -
+Ours - OSTrack-B256 67.5 96.5 61.8 78.6 69.6 83.0
+ViPT (Zhu et al., 2023a) CVPR’23 DropTrack-B384 65.6 93.9 60.0 77.4 68.6 81.4
+SDSTrack (Hou et al., 2024) CVPR’24 DropTrack-B384 65.6 94.4 60.6 78.0 69.0 82.2
+UnTrack (Wu et al., 2024b) CVPR’24 DropTrack-B384 - - 60.0 77.2 - -
+Ours - DropTrack-B384 68.7 97.1 63.0 80.3 71.3 84.9
+MamTrack (Sun et al., 2025) CVPR’25 HiViT-B384 66.4 94.2 61.6 79.2 - -
+SUTrack (Chen et al., 2025b) AAAI’25 SUTrack-B384 - - 63.1 79.6 - -
+Ours - SUTrack-B384 70.8 98.2 65.4 81.5 73.5 85.6
+4.1.2 Evaluation Metrics
+Adhering to recognized standards (Hou et al., 2024; Wu et al., 2024b; Hong et al., 2024; Sun et al., 2025), we evaluated tracking performance with the following metrics.
 For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR).
@@
 Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack.
-For all metrics, the larger, the better.
-IV-A3 Pre-trained Models and Compared Methods
-In this paper, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack [58], the most widely adopted model; DropTrack [49] and SUTrack [9], which offer improved generalization.
-Notably, OSTrack and DropTrack are built upon ViT-B/16 [12], while SUTrack adopts the HiViT-B/16 [64] as the backbone.
-Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack, and SUTrack ($\times192$$\times384$) for template and search regions, respectively.
-To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table I and Table II.
+For all tracking metrics, the larger, the better.
+4.1.3 Pre-trained Models and Compared Methods
+In this work, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack (Ye et al., 2022), the most widely adopted model; DropTrack (Wu et al., 2023) and SUTrack (Chen et al., 2025b), which offer improved generalization and performance.
+Notably, OSTrack and DropTrack are built upon ViT-B/16 (Dosovitskiy et al., 2020), while SUTrack adopts the HiViT-B/16 (Zhang et al., 2023b) as the backbone.
+Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack ($\times192$$\times384$), and SUTrack ($\times192$$\times384$) for template and search regions, respectively.
+To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table 1 and Table 2.
 First, we construct strong RGB-based single-modal baselines under full fine-tuning.
@@
 For fair comparison, all methods are grouped according to the pre-trained trackers.
-Notably, OneTracker [16] and MamTrack [41] adopt non-standard experimental settings and lack open-source implementations.
-OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table III.
+Notably, XTrack (Tan et al., 2025b), OneTracker (Hong et al., 2024) and MamTrack (Sun et al., 2025) adopt non-standard experimental settings and lack open-source implementations.
+XTrack (Tan et al., 2025b) and OneTracker use a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table 3.
 MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
-IV-A4 Training Details
-We follow the data processing pipeline of SDSTrack [17] across all datasets, converting event data into color-polar event images, with no preprocessing for other auxiliary modalities.
+4.1.4 Training Details
+We follow the data processing pipeline of SDSTrack (Hou et al., 2024) across all datasets, converting event data into color-polar event images, with no preprocessing for other auxiliary modalities.
 The models are trained on 8 NVIDIA 3090Ti GPUs with a batch size of 192 and 30 epochs.
@@
 We utilize the AdamW optimizer with a learning rate set to $1\times 10^{-4}$ and a weight decay set to $10^{-4}$.
-TABLE I: Quantitative comparison on the RGB-Event datasets. The best are marked with “bold”, and the second best are marked with “underline”.
-Method Reference Base Model FE108 VisEvent CoeSot
-SR PR SR PR SR PR
-Image-based Methods (Only-RGB)
-DiMP [3] ICCV’19 ResNet 48.5 71.7 45.2 66.2 56.7 69.1
-Stark-S [55] ICCV’21 ResNet 51.4 75.1 44.6 61.3 55.7 66.7
-OSTrack [58] ECCV’22 OSTrack-256 50.7 73.7 53.4 69.6 64.5 77.0
-DropTrack [49] CVPR’23 DropTrack-384 52.0 75.9 54.8 72.4 66.1 79.1
-Cross-modal Transfer Learning
-FENet [61] ICCV’21 DiMP 63.3 92.4 50.8 70.4 53.8 66.1
-AFNet [60] CVPR’23 DiMP 63.8 92.6 51.8 70.8 52.5 63.8
-CEUTrack [42] Arxiv’22 OSTrack-B256 56.7 84.2 55.0 71.6 62.0 76.4
-ProTrack [57] ACM MM’22 OSTrack-B256 58.0 85.5 47.1 63.2 66.1 78.4
-HRTrack [73] ICCV’23 OSTrack-B256 - - - - 63.2 71.9
-ViPT [70] CVPR’23 OSTrack-B256 65.1 93.5 59.2 75.8 67.7 80.6
-SDSTrack [17] CVPR’24 OSTrack-B256 60.0 87.2 59.7 76.7 67.9 80.9
-UnTrack [51] CVPR’24 OSTrack-B256 - - 58.9 75.6 - -
-Ours - OSTrack-B256 67.5 96.5 61.8 78.6 69.6 83.0
-ViPT [70] CVPR’23 DropTrack-B384 65.6 93.9 60.0 77.4 68.6 81.4
-SDSTrack [17] CVPR’24 DropTrack-B384 65.6 94.4 60.6 78.0 69.0 82.2
-UnTrack [51] CVPR’24 DropTrack-B384 - - 60.0 77.2 - -
-Ours - DropTrack-B384 68.7 97.1 63.0 80.3 71.3 84.9
-MamTrack [41] CVPR’25 HiViT-B384 66.4 94.2 61.6 79.2 - -
-SUTrack [9] AAAI’25 SUTrack-B384 - - 63.1 79.6 - -
-Ours - SUTrack-B384 70.8 98.2 65.4 81.5 73.5 85.6
-TABLE II: Quantitative comparison on the RGB-Depth and RGB-Thermal datasets. The best are marked with “bold”, and the second best are marked with “underline”.
-Method Reference Base Model DepthTrack VOT RGBD2022 LasHeR RGBT234
-Pr Re F-score Acc Rob EAO SR PR MSR MPR
-Image-based Methods (Only-RGB)
-DiMP [3] ICCV’19 ResNet 46.3 42.8 44.5 70.3 73.1 54.3 42.8 54.4 42.1 62.5
-Stark-S [55] ICCV’21 ResNet 39.3 37.6 38.4 65.4 62.8 48.2 37.7 46.3 48.9 66.5
-OSTrack [58] ECCV’22 OSTrack-B256 53.6 52.2 52.9 80.3 83.3 67.6 47.2 58.5 54.9 72.9
-DropTrack [49] CVPR’23 DropTrack-B384 56.4 55.8 56.1 81.5 85.1 69.2 47.7 59.4 57.2 75.8
-Cross-modal Transfer Learning
-SPT [71] AAAI’23 Stark-S 52.7 54.9 53.8 79.8 85.1 65.1 33.4 46.7 55.5 78.6
-APFNet [53] AAAI’22 Stark-S 51.6 51.4 51.5 79.0 83.7 64.2 36.2 50.0 57.9 82.7
-ProTrack [57] ACM MM’22 OSTrack-B256 58.3 57.3 57.8 80.1 80.2 65.1 42.0 53.8 59.9 79.5
-ViPT [70] CVPR’23 OSTrack-B256 59.2 59.6 59.4 81.5 87.1 72.1 52.5 65.1 61.7 83.5
-SDSTrack [17] CVPR’24 OSTrack-B256 61.9 60.9 61.4 81.2 88.3 72.8 53.1 66.5 62.5 84.8
-UnTrack [51] CVPR’24 OSTrack-B256 61.1 60.8 61.0 82.0 86.9 72.1 51.3 63.7 62.5 84.2
-Ours - OSTrack-B256 64.7 65.4 65.1 82.1 88.8 74.1 56.3 70.1 64.4 87.2
-ViPT [70] CVPR’23 DropTrack-B384 62.6 61.9 62.2 81.7 87.3 72.3 52.5 65.6 63.4 84.7
-SDSTrack [17] CVPR’24 DropTrack-B384 63.3 62.2 62.7 81.4 88.6 73.0 55.1 69.0 65.0 87.1
-UnTrack [51] CVPR’24 DropTrack-B384 62.4 61.7 62.0 82.1 87.1 72.2 52.4 65.9 64.1 85.2
-Ours - DropTrack-B384 67.2 67.1 67.1 82.9 89.2 74.5 58.8 73.6 66.7 89.1
-MamTrack [41] CVPR’25 HiViT-B384 - - - - - - 54.2 67.4 62.4 84.4
-SUTrack [9] AAAI’25 SUTrack-B384 58.8 58.6 58.7 83.0 90.6 75.4 60.8 75.6 69.2 92.1
-Ours - SUTrack-B384 65.1 65.2 65.2 84.1 92.6 77.7 62.9 77.8 70.3 93.3
-TABLE III: Quantitative comparison with OneTracker.
-Method Reference Base Model VisEvent DepthTrack VOT RGBD2022 LasHeR RGBT234
-SR PR Pr Re F-score Acc Rob EAO SR PR MSR MPR
-OneTracker [16] CVPR’24 OSTrack-B384 60.8 76.7 60.7 60.4 60.9 81.9 87.2 72.7 53.8 67.2 64.2 85.7
-Ours - OSTrack-B384 62.6 79.4 65.3 66.1 65.6 82.4 88.9 74.3 57.0 71.2 65.4 87.9
-IV-B Comparison with State-of-the-Art Methods
-Extensive comparative analyses are presented in Table I, Table II and Table III, where our method demonstrates excellent performance on all multi-modal benchmarks after incorporating the proposed regularized tuning strategies.
+4.2 Comparison with State-of-the-Art Methods
+Extensive comparative analyses are presented in Table 1, Table 2 and Table 3, where our method demonstrates excellent performance on all multi-modal tracking benchmarks after incorporating the proposed regularized tuning strategies.
 The corresponding precision and success plots are illustrated in Fig.
@@
 Crucially, the remarkable improvements achieved by our method suggest the significance and necessity of developing tailored cross-domain transfer techniques for multi-modal object tracking.
+Table 2: Quantitative comparison on the RGB-Depth and RGB-Thermal datasets. The best are marked with “bold”, and the second best are marked with “underline”.
+Method Reference Base Model DepthTrack VOT RGBD2022 LasHeR RGBT234
+Pr Re F-score Acc Rob EAO SR PR MSR MPR
+Image-based Methods (Only-RGB)
+DiMP (Bhat et al., 2019) ICCV’19 ResNet 46.3 42.8 44.5 70.3 73.1 54.3 42.8 54.4 42.1 62.5
+Stark-S (Yan et al., 2021a) ICCV’21 ResNet 39.3 37.6 38.4 65.4 62.8 48.2 37.7 46.3 48.9 66.5
+OSTrack (Ye et al., 2022) ECCV’22 OSTrack-B256 53.6 52.2 52.9 80.3 83.3 67.6 47.2 58.5 54.9 72.9
+DropTrack (Wu et al., 2023) CVPR’23 DropTrack-B384 56.4 55.8 56.1 81.5 85.1 69.2 47.7 59.4 57.2 75.8
+Cross-modal Transfer Learning
+SPT (Zhu et al., 2023b) AAAI’23 Stark-S 52.7 54.9 53.8 79.8 85.1 65.1 33.4 46.7 55.5 78.6
+APFNet (Xiao et al., 2022) AAAI’22 Stark-S 51.6 51.4 51.5 79.0 83.7 64.2 36.2 50.0 57.9 82.7
+ProTrack (Yang et al., 2022) ACM MM’22 OSTrack-B256 58.3 57.3 57.8 80.1 80.2 65.1 42.0 53.8 59.9 79.5
+ViPT (Zhu et al., 2023a) CVPR’23 OSTrack-B256 59.2 59.6 59.4 81.5 87.1 72.1 52.5 65.1 61.7 83.5
+SDSTrack (Hou et al., 2024) CVPR’24 OSTrack-B256 61.9 60.9 61.4 81.2 88.3 72.8 53.1 66.5 62.5 84.8
+UnTrack (Wu et al., 2024b) CVPR’24 OSTrack-B256 61.1 60.8 61.0 82.0 86.9 72.1 51.3 63.7 62.5 84.2
+Ours - OSTrack-B256 64.7 65.4 65.1 82.1 88.8 74.1 56.3 70.1 64.4 87.2
+ViPT (Zhu et al., 2023a) CVPR’23 DropTrack-B384 62.6 61.9 62.2 81.7 87.3 72.3 52.5 65.6 63.4 84.7
+SDSTrack (Hou et al., 2024) CVPR’24 DropTrack-B384 63.3 62.2 62.7 81.4 88.6 73.0 55.1 69.0 65.0 87.1
+UnTrack (Wu et al., 2024b) CVPR’24 DropTrack-B384 62.4 61.7 62.0 82.1 87.1 72.2 52.4 65.9 64.1 85.2
+Ours - DropTrack-B384 67.2 67.1 67.1 82.9 89.2 74.5 58.8 73.6 66.7 89.1
+MamTrack (Sun et al., 2025) CVPR’25 HiViT-B384 - - - - - - 54.2 67.4 62.4 84.4
+SUTrack (Chen et al., 2025b) AAAI’25 SUTrack-B384 58.8 58.6 58.7 83.0 90.6 75.4 60.8 75.6 69.2 92.1
+Ours - SUTrack-B384 65.1 65.2 65.2 84.1 92.6 77.7 62.9 77.8 70.3 93.3
+Table 3: Quantitative comparison with OneTracker and XTrack.
+Method Reference Base Model VisEvent DepthTrack VOT RGBD2022 LasHeR RGBT234
+SR PR Pr Re F-score Acc Rob EAO SR PR MSR MPR
+OneTracker (Hong et al., 2024) CVPR’24 OSTrack-B384 60.8 76.7 60.7 60.4 60.9 81.9 87.2 72.7 53.8 67.2 64.2 85.7
+XTrack (Tan et al., 2025b) ICCV’25 OSTrack-B384 60.9 77.5 61.8 62.0 61.5 82.1 88.8 74.0 55.7 69.1 64.9 87.4
+Ours - OSTrack-B384 62.6 79.4 65.3 66.1 65.6 82.4 88.9 74.3 57.0 71.2 65.4 87.9
 Results on RGB-Event Tracking.
-As shown in Table I and Table III, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively.
+As shown in Table 1 and Table 3, our method surpasses all state-of-the-art trackers across all RGB-Event datasets, achieving the highest precision scores of $98.2\%$, $81.5\%$ and $85.6\%$ on the FE108, VisEvent, and CoeSot datasets, respectively.
 In particular, on FE108, a dataset characterized by low-light conditions and heavy reliance on event information, our approach surpasses the previous best by a notable margin: $3.0\%$ in PR and $2.4\%$ in SR over OSTrack-B256.
@@
 Results on RGB-Depth Tracking.
-As depicted in Table II and Table III, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score.
+As depicted in Table 2 and Table 3, our method outperforms all previous state-of-the-art trackers on DepthTrack, obtaining the top performance $67.1\%$ in F-score.
 Using the pre-trained OSTrack-B256, our method yields substantial improvements: $2.8\%$ in Pr, $4.5\%$ in Re, and $3.7\%$ in F-score.
@@
 Furthermore, despite no training on the VOT-RGBD2022 dataset, our method still delivers enhanced performance with a $2.3\%$ gain in EAO over existing baselines.
-This clearly demonstrates the strong generalization and transferability of our approach in cross-modal tracking scenarios.
+This clearly demonstrates the strong transferability of our approach in cross-modal tracking scenarios.
 Results on RGB-Thermal Tracking.
-As listed in Table II and Table III, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256.
+As listed in Table 2 and Table 3, our method achieves $70.1\%$ precision and $56.3\%$ success when evaluated on the pre-trained OSTrack-B256.
 Furthermore, our method effectively unleashes the potential of the DropTrack, yielding substantial improvements of $4.6\%$ in PR, $3.7\%$ in SR.
@@
 Moreover, our method also maintains superior performance across other challenging attributes such as Out-of-View and Frame Lost.
-IV-C Ablation Study and Analysis
-We conducted ablative analyses to verify the effectiveness and characteristics of sensitivity-aware regularization tuning.
-In the ablation section, “full fine-tuning” refers to training the entire backbone, excluding the box head (as Section III-A).
-All methods use the pre-trained OSTrack-B256 weights unless otherwise specified.
-Effectiveness of Proposed Sensitivities.
-We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Table IV.
-A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability.
-Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior sensitivity regularization scheme significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness.
-Similarly, the comparison between (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements.
-Notably, the simultaneous application of both prior and transfer sensitivity yields significantly greater improvements than either method alone, proving their complementary nature.
-While the transfer sensitivity penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
-TABLE IV: Ablative study results of the proposed key components.
-“FFT” refers to full fine-tuning of the backbone; “PS” represents prior sensitivity; “TS” is transfer sensitivity.
+Overall, our models performs substantially better across diverse challenging conditions.
+4.3 Ablation Study and Analysis
+We conducted ablative analyses to verify the effectiveness and characteristics of significance-aware regularization tuning.
+In the ablation section, “full fine-tuning” refers to training the entire backbone, excluding the box head (as in Section 3.1).
+All methods use the pre-trained OSTrack-B256 weights unless specified otherwise.
+Effectiveness of Parameter Significance.
+We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed parameter significance items, as summarized in Table 4.
+Comparing (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability.
+Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior significance regularization significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness.
+Similarly, comparing (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements.
+Notably, the simultaneous application of both prior and transfer significance yields greater improvements than either method alone, proving their complementary nature.
+While the transfer significance penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
+Table 4: Ablative study results of the proposed key components.
+“FFT” refers to full fine-tuning of the backbone; “PS” represents prior significance; “TS” is transfer significance.
 (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
@@
 (e) $\checkmark$ $\checkmark$ $\checkmark$ 67.4 96.5 64.7 65.4 65.1 56.3 70.1
-Figure 8: Operation-wise eigenvalue spectra for measuring prior sensitivity.
-Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale.
-For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
-Eigenvalue Budget (Top-$K$) for Prior Sensitivity.
-In this section, we examine the effect of the eigenvalue count $K$ in Eq.
-(8).
-For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them.
-As illustrated in Fig.
-8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues.
-Consistent with this observation, Table V shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline.
-This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes.
-Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
-TABLE V: Ablation analysis of low-dimension eigenvalue count $K$ on prior sensitivity estimation.
+Table 5: Ablation analysis of low-dimension eigenvalue count $K$ on prior significance estimation.
 Exp. FE108 DepthTrack LasHeR
@@
 KK=20 66.4 95.2 64.1 64.4 64.2 55.4 68.9
-Sensitivity Harmony Coefficient $\kappa$.
-In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq.
-(14), with the detailed results presented in Table VI.
+Figure 8: Operation-wise eigenvalue spectra for measuring prior significance.
+Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale.
+For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
+Eigenvalue Budget (Top-$K$) for Prior Significance.
+In this section, we examine the effect of the eigenvalue count $K$ in Eq.
+(8).
+For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior significance using the top-$K$ of them.
+As illustrated in Fig.
+8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues.
+Consistent with this observation, Table 5 shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline.
+This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of significance rather than their absolute magnitudes.
+Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
+Significance Harmony Coefficient $\kappa$.
+In this section, we evaluate the impact of different significance harmony coefficients, $\kappa$ as defined in Eq.
+(14), with the detailed results presented in Table 6.
 By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9).
-This finding suggests that while both prior sensitivity and transfer sensitivity contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
-TABLE VI: Ablation analysis of sensitivity harmony coefficient on the FE108 dataset.
+This finding suggests that while both prior significance and transfer significance contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
+Additional significance fusion scheduling strategies are provided in Appendix C.
+Table 6: Ablation analysis of significance harmony coefficient on the FE108 dataset.
 $\kappa$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
@@
 To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well.
-As shown in Table VII, unfreezing the box head further exacerbates overfitting.
+As shown in Table 7, unfreezing the box head further exacerbates overfitting.
 While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge.
-TABLE VII: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the sensitivity-aware regularization tuning method.
+Table 7: Ablation results of box head tuning. “Box” denotes tuning the box head, “S-Reg” represents the significance-aware regularization tuning method.
 Exp. Box FE108 DepthTrack LasHeR
@@
 Compatibility with PEFT Methods.
-Existing PEFT methods [70, 17, 51] typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization.
+Existing PEFT methods Zhu et al.
+(2023a); Hou et al.
+(2024); Wu et al.
+(2024b) typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization.
 To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied.
-As shown in Table VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR.
+As shown in Table 8, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR.
 Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR.
@@
 However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
-TABLE VIII: Compatibility study results of the proposed regularization tuning method on existing PEFT methods.
-“FFT” denotes full fine-tuning of the backbone, while “S-Reg” represents the sensitivity-aware regularization tuning scheme.
+Table 8: Compatibility study results of the proposed regularization tuning method on existing PEFT methods.
+“FFT” denotes full fine-tuning of the backbone, while “S-Reg” represents the significance-aware regularization tuning scheme.
 Exp. VisEvent DepthTrack LasHeR
@@
 To investigate this, we conduct ablation studies highlighting their impact on different modalities.
-As shown in Table IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques.
+As shown in Table 9, both RGB and auxiliary modalities benefit significantly from our regularization techniques.
 For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset.
@@
 These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
-TABLE IX: Compatibility study results of the proposed regularization tuning method on single-modal data. “S-Reg” represents the regularization tuning method.
+Table 9: Compatibility study results of the proposed regularization tuning method on single-modal data. “S-Reg” represents the regularization tuning method.
 Exp. CoeSot DepthTrack LasHeR
@@
 The auxiliary branches are derived from the VisEvent, DepthTrack, and LasHeR datasets, respectively.
-Impact of Event Representations.
+Impact of Input Event Representations.
 In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer.
 In this experiment, we keep the same training setup, varying only the input event representations.
-As shown in Table X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps.
+As shown in Table 10, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps.
 In contrast, count-based representations (Event Count, Event Frame) offer more robust performance.
@@
 Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
-TABLE X: Ablation results of different event representations.
+Table 10: Ablation results of different event representations.
 Exp. FE108 VisEvent CoeSot
 SR PR SR PR SR PR
-Event Frame [38] 66.8 95.5 61.2 77.6 68.8 82.0
-Event Count [33] 66.3 95.1 59.8 76.9 67.9 81.2
-Time Surface [39] 66.1 94.9 60.3 77.1 67.4 80.6
-TSLTD [6] 66.5 95.2 61.0 77.4 67.9 81.3
-Event Volume [69] 67.3 96.1 61.4 77.9 69.1 82.4
+Event Frame Rebecq et al. (2017) 66.8 95.5 61.2 77.6 68.8 82.0
+Event Count Maqueda et al. (2018) 66.3 95.1 59.8 76.9 67.9 81.2
+Time Surface Sironi et al. (2018) 66.1 94.9 60.3 77.1 67.4 80.6
+TSLTD Chen et al. (2020) 66.5 95.2 61.0 77.4 67.9 81.3
+Event Volume Zhu et al. (2019) 67.3 96.1 61.4 77.9 69.1 82.4
 Color-Polar Event Image (Ours) 67.5 96.5 61.8 78.6 69.6 83.0
-Effectiveness of Smaller Learning Rates.
-In this paper, we exploit the sensitivity-regularized gradients to bolster the transfer process.
-A straightforward strategy might involve using a smaller learning rate.
-To investigate this, we evaluate the effect of reduced learning rates.
-As shown in Table XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR.
-Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation.
-These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
-TABLE XI: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
+Table 11: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
 Exp. $\alpha=10^{-4}$ $\alpha=10^{-5}$ $\alpha=10^{-6}$ Ours
@@
 PR 65.8 66.2 65.4 70.1
+Effectiveness of Smaller Learning Rates.
+In this paper, we exploit the significance-regularized gradients to bolster the transfer process.
+A straightforward strategy might involve using a smaller learning rate.
+To investigate this, we evaluate the effect of reduced learning rates.
+As shown in Table 11, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR.
+Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation.
+These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
+Figure 10: Significance-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
 Observations on Weight Variations.
-To intuitively demonstrate the effectiveness of the proposed sensitivity-aware regularization tuning, we visualize parameter dynamics during training from both spatial and temporal perspectives, as depicted in Fig.
+To intuitively demonstrate the effectiveness of the proposed significance-aware regularization tuning, we visualize parameter dynamics during training from both spatial and temporal perspectives, as depicted in Fig.
 9 and Fig.
@@
 As shown in Fig.
-9, for parameters with high prior sensitivity, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates.
+9, for parameters with high prior significance, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates.
 As further demonstrated in Fig.
-10, our method facilitates faster convergence and achieves a substantial reduction in sensitivity-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation.
+10, our method facilitates faster convergence and achieves a substantial reduction in significance-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation.
 Moreover, the vanilla fine-tuning causes parameter updates to concentrate excessively in localized regions, hindering the global adaptability.
-After applying the sensitivity-aware penalties, the parameter updates become more balanced and distributed, effectively decentralizing the adaptation burden from a few parameters to a broader set.
+After applying the significance-aware penalties, the parameter updates become more balanced and distributed, effectively decentralizing the adaptation burden from a few parameters to a broader set.
 This redistribution helps mitigate adaptation risk and promotes more stable fine-tuning.
-Figure 10: Sensitivity-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
-TABLE XII: Computational cost of Prior Sensitivity Estimation on OSTrack-B256.
-Metric Computational Cost
-Operation-wise Model-wise Total Process
-FLOPs 104.8 G 15.5 T $\sim$ 39.7 E
-Latency $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
-Memory $164.2$ GB - -
-TABLE XIII: Computational complexity and speed analysis on the LasHeR dataset.
-Method Base Model Param (M) FLOPs (G) FPS SR PR
-OSTrack [58] OSTrack 92.1 58.1 98.4 47.2 58.5
-ProTrack [57] OSTrack 92.7 58.4 92.3 42.0 53.8
-ViPT [70] OSTrack 92.9 59.9 88.6 52.5 65.1
-SDSTrack [17] OSTrack 102.1 108.7 44.6 53.1 66.5
-UnTrack [51] OSTrack 98.7 62.6 75.6 51.3 63.7
-ViPT+Ours OSTrack 92.9 59.9 88.6 54.9 68.4
-UnTrack+Ours OSTrack 98.7 62.6 75.6 54.0 67.4
-Ours OSTrack 202.0 149.0 49.0 56.3 70.1
-DropTrack [49] DropTrack 92.1 130.6 57.6 47.7 59.2
-ViPT [70] DropTrack 92.9 131.9 48.8 52.5 65.6
-SDSTrack [17] DropTrack 102.1 244.2 26.8 55.1 69.0
-UnTrack [51] DropTrack 98.7 138.2 42.3 52.4 65.9
-ViPT+Ours DropTrack 92.9 131.9 48.8 57.1 70.5
-UnTrack+Ours DropTrack 98.7 138.2 42.3 56.3 69.4
-Ours DropTrack 202.0 335.3 29.2 58.8 73.6
-SUTrack [9] SUTrack 65.29 106.9 37.6 60.8 75.6
-Ours SUTrack 141.1 243.7 24.8 62.9 77.8
-IV-D Computational Efficiency
-In addition to tracking accuracy, we evaluate the computational efficiency of our sensitivity-aware regularization tuning by analyzing both training and inference overhead.
-Computational Overhead of Prior Sensitivity.
-Prior sensitivity is estimated via a group-wise FIM approximation with eigen-decomposition.
-With the operation-diagonal (group-wise) approximation and Rayleigh-quotient probing, we reduce the computational cost from intractable to a practical level.
-We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table XIV to quantify its efficiency.
-While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior sensitivity estimation introduces additional, non-trivial computation.
-Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference.
-Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle (see Appendix for details).
-Efficiency of Sensitivity-Regularized Fine-tuning.
-We compare the training efficiency of our method against vanilla fine-tuning.
-Since prior sensitivity estimation is conducted as a one-time preprocessing step before fine-tuning, its runtime is excluded from the training time.
-Our method introduces parameter-wise transfer sensitivity quantification with marginal computational overhead, as this sensitivity is derived from the off-the-shelf gradients.
-While vanilla fine-tuning operates at 37.5 ms per iteration, our method runs at 39.0 ms per iteration, incurring only a modest $4.0\%$ increase in training time.
+Table 12: Computational complexity and speed analysis on the LasHeR dataset. The base trackers are denoted by abbreviations (e.g., OS/Drop/SU). Model size is reported in millions of parameters (M).
+Method Model Params GFLOPs FPS SR PR
+OSTrack Ye et al. (2022) OS 92.1 58.1 98.4 47.2 58.5
+ProTrack Yang et al. (2022) OS 92.7 58.4 92.3 42.0 53.8
+ViPT Zhu et al. (2023a) OS 92.9 59.9 88.6 52.5 65.1
+SDSTrack Hou et al. (2024) OS 102.1 108.7 44.6 53.1 66.5
+UnTrack Wu et al. (2024b) OS 98.7 62.6 75.6 51.3 63.7
+ViPT+Ours OS 92.9 59.9 88.6 54.9 68.4
+UnTrack+Ours OS 98.7 62.6 75.6 54.0 67.4
+Ours OS 202.0 149.0 49.0 56.3 70.1
+DropTrack Wu et al. (2023) Drop 92.1 130.6 57.6 47.7 59.2
+ViPT Zhu et al. (2023a) Drop 92.9 131.9 48.8 52.5 65.6
+SDSTrack Hou et al. (2024) Drop 102.1 244.2 26.8 55.1 69.0
+UnTrack Wu et al. (2024b) Drop 98.7 138.2 42.3 52.4 65.9
+ViPT+Ours Drop 92.9 131.9 48.8 57.1 70.5
+UnTrack+Ours Drop 98.7 138.2 42.3 56.3 69.4
+Ours Drop 202.0 335.3 29.2 58.8 73.6
+SUTrack Chen et al. (2025b) SU 65.29 106.9 37.6 60.8 75.6
+Ours SU 141.1 243.7 24.8 62.9 77.8
+4.4 Computational Efficiency
+In addition to tracking accuracy, we evaluate the computational efficiency of our significance-aware regularization tuning by analyzing both training and inference overhead.
+Computational Overhead of Prior Significance.
+Our method requires an offline estimation of prior significance (e.g., FIM-based eigen-decomposition and Rayleigh quotient probing), which introduces a non-trivial preprocessing cost compared to purely plug-and-play PEFT baselines.
+This overhead can become a practical barrier when scaling to larger backbones or when frequent architectural changes demand repeated recomputation.
+Nonetheless, this cost is paid once and does not materially affect the per-iteration training and inference runtime after the significance is computed (details in the Appendix C).
+Efficiency of Significance-Regularized Fine-tuning.
+We compare the training efficiency of our SRFT against vanilla FFT.
+Since prior significance is computed once as a preprocessing step, we exclude it from per-iteration training time.
+SRFT introduces parameter-wise transfer significance quantification with marginal computational overhead, as this significance is derived from the off-the-shelf gradients: 39.0 ms/iter vs.
+37.5 ms/iter for vanilla FFT ($4.0\%$ ).
 Inference Cost and Speed.
@@
 All experiments are conducted on the same hardware platform, equipped with an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and a single NVIDIA RTX 3090Ti GPU.
-As reported in Table XIII, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting.
+As reported in Table 12, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting.
 When integrated into ViPT (“ViPT+Ours”), it yields both superior speed and performance.
 Under the DropTrack base model, the resulting tracker still operates at 48.8 FPS, maintaining a respectable speed given its focus on accuracy.
-IV-E Limitations and Discussion
-From a computational-efficiency perspective, a practical consideration is that prior sensitivity estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles.
+4.5 Limitations and Discussion
+From a computational-efficiency perspective, a practical consideration is that prior significance estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles.
 Additionally, the formulation defines the Fisher expectation over the pre-training dataset, so applying the method as-is may require access to (or a governed surrogate for) pre-training data—an assumption that does not always hold under licensing, privacy, or retention constraints—potentially limiting portability across model sources and deployment settings.
-Addressing these constraints, e.g., by exploring data-free or proxy-based sensitivity estimation—remains an important direction for future work.
-V Conclusion
+Addressing these constraints, e.g., by exploring data-free or proxy-based significance estimation—remains an important direction for future work.
+5 Conclusion
 This study revisits the critical misfitting issues encountered when adapting pre-trained RGB trackers to multi-modal tracking tasks.
-By introducing two complementary sensitivities that capture the dynamic shift in parameter importance as models transition from pre-trained to multi-modal contexts, our regularized tuning method strategically calibrates gradient updates.
+By introducing two complementary parameter significance that capture the dynamic shift in parameter importance as models transition from pre-trained to multi-modal contexts, our regularized tuning method strategically calibrates gradient updates.
 Extensive experiments demonstrate the effectiveness of our method, surpassing current state-of-the-art techniques across various multi-modal tracking scenarios, with notable performance improvements resulting from the incorporation of regularization.
-A key insight from our work is the importance of sensitivity-aware fine-tuning, which offers superior generalization and adaptability compared to traditional fine-tuning methods.
+A key insight from our work is the importance of significance-aware fine-tuning, which offers superior generalization and adaptability compared to traditional fine-tuning methods.
 Ultimately, our findings highlight the need for more nuanced cross-domain generalization strategies, as overly rigid or flexible fine-tuning can hinder pre-trained model transferability.
 We believe these insights will pave the way for further advancements in multi-modal transfer learning, particularly in the context of scene perception tasks.
-Appendix A Proof of Derivations
-A-A Proof of Proposition 1
-  1.
+Data Availability
+This work does not propose a new dataset.
+All the datasets we used are publicly available and well cited in the paper.
+FE108: https://zhangjiqing.com/dataset/; VisEvent: https://sites.google.com/view/viseventtrack/; CoeSot: https://github.com/Event-AHU/COESOT; DepthTrack: https://github.com/xiaozai/det; VOT RGBD2022: https://www.votchallenge.net/vot2022/dataset.html; LasHeR: https://github.com/BUGPLEASEOUT/LasHeR; RGBT234: https://github.com/xuboyue1999/RGBT-Tracking/tree/main.
+Conflict of Interest
+The authors affirm that there are no commercial or associative relationships that could be perceived as a conflict of interest related to the submitted work.
+Appendix
+This appendix contains the following contents.
+In Section A, we provide implementation details for multi-modal tracking, including a more detailed description of the network architecture.
+In Section B, we present the complete derivations and proof details.
+In Section C, we report additional quantitative results, including comparisons with SPT, the computational overhead introduced by prior significance estimation, the training efficiency of SRFT, and different significance fusion scheduling strategies.
+Finally, we have supplemented some tracking visuals for a better qualitative comparison in Section D.
+  • Section A: Implementation details.
+  • Section B: Proof of derivations.
+  • Section C: Additional ablation studies.
+  • Section D: Visualization of tracking results.
+Appendix A Implementation details
+The detailed structure of the multimodal tracker is shown in Fig.
+11.
+The input of our proposed tracking model consists of a pair of template frames and a pair of search frames, i.e., one RGB template frame $\mathrm{Z}_{\mathrm{R}}\in\mathbb{R}^{H_{z}\times W_{z}\times 3}$, one RGB search frame $\mathrm{X}_{\mathrm{R}}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, one auxiliary-modal template frame $\mathrm{Z}_{\mathrm{A}}\in\mathbb{R}^{H_{z}\times W_{z}\times 3}$, and one auxiliary-modality search frame $\mathrm{X}_{\mathrm{A}}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$.
+Notably, to make event data compatible with the RGB domain, we aggregate the event set between the image and its next one into a three-channel color-polar event image.
+These data are first split and flattened into sequences of patches $\mathrm{z}_{\mathrm{R}},\mathrm{z}_{\mathrm{A}}\in\mathbb{R}^{N_{z}\times\left(3P^{2}\right)}$ and $\mathrm{x}_{\mathrm{R}},\mathrm{x}_{\mathrm{A}}\in\mathbb{R}^{N_{x}\times\left(3P^{2}\right)}$, where $P\times P$ is the resolution of each patch ($P=16$), and $N_{z}=\frac{H_{z}W_{z}}{P^{2}}$, $N_{x}=\frac{H_{x}W_{x}}{P^{2}}$.
+Next, two modal-aware patch embedding layers are used to project $\mathrm{z}_{\mathrm{R}},\mathrm{x}_{\mathrm{R}}$ and $\mathrm{z}_{\mathrm{A}},\mathrm{x}_{\mathrm{A}}$ into the D-dimensional latent space, $\mathrm{z}_{\mathrm{R}},\mathrm{z}_{\mathrm{A}}\in\mathbb{R}^{N_{z}\times D}$ and $\mathrm{x}_{\mathrm{R}},\mathrm{x}_{\mathrm{A}}\in\mathbb{R}^{N_{x}\times D}$.
+The patch embeddings $\mathrm{z}_{\mathrm{R}}$ and $\mathrm{x}_{\mathrm{R}}$ are concatenated as $\mathbf{H}_{\mathrm{R}}^{(0)}=\left[\mathrm{z}_{\mathrm{R}};\mathrm{x}_{\mathrm{R}}\right]\in\mathbb{R}^{\left(N_{z}+N_{x}\right)\times D}$, and $\mathrm{z}_{\mathrm{A}}$ and $\mathrm{x}_{\mathrm{A}}$ are concatenated as $\mathbf{H}_{\mathrm{A}}^{(0)}=\left[\mathrm{z}_{\mathrm{A}};\mathrm{x}_{\mathrm{A}}\right]\in\mathbb{R}^{\left(N_{z}+N_{x}\right)\times D}$.
+The computation of modality-aware ViT block can be formulated as:
+$\begin{gathered}\mathbf{H}_{\mathrm{X}}^{\prime(l)}=\mathbf{H}_{\mathrm{X}}^{(l-1)}+\operatorname{MSA}\left(\mathrm{LN}\left(\mathbf{H}_{\mathrm{X}}^{(l-1)}\right)\right)\\ \mathbf{H}_{\mathrm{X}}^{(l)}=\mathbf{H}_{\mathrm{X}}^{\prime(l)}+\operatorname{MLP}\left(\mathrm{LN}\left(\mathbf{H}_{\mathrm{X}}^{\prime(l)}\right)\right)\end{gathered}$
+[…]
+Figure 11: More detailed network architecture of our multi-modality trackers.
+Appendix B Proof of Derivations
+B.1 Validity of Loss-Parameter Manifold Hypothesis
+The loss-parameter manifold hypothesis is the mathematical foundation of modern deep learning optimization.
+It posits that, for a given model and dataset, the loss function induces a highly structured geometric object—a manifold (or manifold-like set)—embedded in parameter space, and that training dynamics can be understood as motion on or near this surface.
+For the manifold hypothesis, a large body of prior work (Bengio et al., 2013; Pope et al., 2021; Song and Ermon, 2019; Kiani et al., 2024) has widely and successfully employed manifold-based assumptions to analyze representation learning and generalization behavior in deep networks.
+While these works consistently demonstrate that the manifold assumption may not hold globally, local manifold structure provides a meaningful and empirically validated abstraction for reasoning about model behavior under realistic perturbations—precisely the regime considered in our analysis.
+In this view, parameters with similar loss values form smooth level sets, locally approximable by tangent planes, while curvature governs how sensitive the loss is to perturbations in different directions.
+In this work, the prior significance does not assume that a global manifold hypothesis holds in multi-modality tracking scenarios.
+Instead, our method is deliberately local: it characterizes the behavior of the pretrained objective in a neighborhood of the pretrained parameters $\theta_{0}$ with pre-trained dataset $D_{0}$, where our prior significance estimation operates.
+As discussed in Section 3, the prior significance approximation defines a practical and computable measure of the increase in pretrained loss induced by parameter deviations around $\theta_{0}$ under $D_{0}$, where the model gradient vanishes; this increase is characterized by the second-order Fisher information matrix (FIM).
+This formulation serves as a geometric metric, identifying the directions in parameter space most crucial for preserving pretrained knowledge, rather than making a claim about global parameter geometry.
+To ensure practical applicability, we leverage the empirically observed spectral concentration of the FIM in deep networks and adopt a low-rank approximation via Rayleigh-quotient probing, supported by an approximation error analysis.
+B.2 Proof of Proposition 1
+  1. 1.
     Bounded FIM Error.
@@
     In particular, under the common assumption that the top-$K$ eigenvalues in each group capture the most significant curvature, where $\mathrm{Tr}(\mathcal{F}^{(\theta_{0})}-\tilde{\mathcal{F}}^{(\theta_{0})})\ll\mathrm{Tr}(\mathcal{F}^{(\theta_{0})})$, the residual eigenvalue mass is small.
-  2.
+  2. 2.
     Bounded Generalization Gap Error.
@@
     In practice, by retaining the top-$K$ eigenvalues in each FIM block (which capture the majority of the FIM energy) and replacing the remaining eigenvalues with an isotropic average, one obtains a low-rank Fisher matrix approximation that preserves the important Riemannian geometry of the parameter space while bounding the distortion in any distance or generalization metric by the residual eigenvalue mass.
-A-B Proof of Transfer Sensitivity
-We state the local assumptions under which the transfer sensitivity bound in Eq. (9)-(12) holds and provide a more explicit derivation.
+B.3 Derivation of Transfer Significance
+We state the local assumptions under which the transfer parameter significance bound in Eq. (9)-(12) holds and provide a more explicit derivation.
 Assumptions.
@@
 Setup.
-Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall
+Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall:
 $\epsilon_{\text{ada}}=\mathbb{E}\left[|S\left(\theta,\delta\mid M\right)-S\left(\theta,\delta^{\prime}\mid M\right)\right|],$ (17)
 $S(\theta,\delta\mid M)=\mathcal{L}(\theta\mid M)-\mathcal{L}(\theta+\delta\mid M).$ (18)
-We model perturbations around a single update step as
+We model perturbations around a single update step as:
 $\delta=\alpha(\mathcal{\mathcal{G}}+\xi),\qquad\delta^{\prime}=\alpha(\mathcal{G}+\xi^{\prime}),\qquad$ (19)
@@
 Assume $L(\theta\mid M)$ is differentiable and locally $\beta$-smooth around $\theta$, i.e., $\|\nabla\mathcal{L}(\theta+\delta\mid M)-\nabla\mathcal{L}(\theta\mid M)\|_{2}\leq\beta\|\delta\|_{2}$.
-Then a first-order Taylor expansion yields
+Then a first-order Taylor expansion yields:
 $\mathcal{L}(\theta+\delta\mid M)=\mathcal{L}(\theta\mid M)+\mathcal{G}^{\top}\delta+r(\delta),\quad|r(\delta)|\leq\frac{\beta}{2}\|\delta\|_{2}^{2},$ (20)
-and hence
-$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta).$ (21)
-Therefore,
-$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})).$ (22)
-When perturbations are sufficiently small (so the remainder term is dominated), the leading term is
-$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi).$ (23)
+and hence:
+$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta),$ (21)
+Therefore, we obtain:
+$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})),$ (22)
+When perturbations are sufficiently small (so the remainder term is dominated), the leading term is:
+$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi),$ (23)
 A guaranteed upper bound.
 Let $X:=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi)$.
-Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus
-$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (24)
-Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized sensitivity:
-$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (25)
+Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus,
+$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (24)
+Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized significance:
+$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (25)
 This shows that $\epsilon_{\text{ada}}$ scales at most linearly with $\|\mathcal{G}\|_{2}$.
@@
 so the linear scaling is tightest in locally smooth regions and for sufficiently small perturbations.
-Appendix B Additional ablation studies
-B-A Computational Overhead of Prior Sensitivity under Larger Backbones
-In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior sensitivity estimation under larger backbones, such as ViT-L and Swin-L.
-Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Tab.
-XIV.
+Appendix C Additional Ablation Studies
+C.1 Comparison with Sensitivity-aware Sparse Tuning (SPT).
+In our method, parameter significance is used to regularize the update of all parameters, rather than to select a subset of “sensitive” parameters for sparse fine-tuning as in SPT (He et al., 2023).
+This distinction is particularly important for cross-modal adaptation.
+To examine this, we compare our approach with SPT under identical training settings, varying only the trainable parameter ratio $\tau$ for SPT (top-$\tau$ sensitive parameters).
+As reported in Tab.
+13, SPT exhibits a clear performance–sparsity trade-off: when $\tau$ decreases, tracking accuracy drops sharply (e.g., F-score $61.0\rightarrow 60.2\rightarrow 54.5$ for $\tau=50\%,20\%,10\%$), indicating that cross-modal adaptation requires distributed parameter updates rather than a highly sparse subset.
+Even with a relatively large budget ($\tau=50\%$), SPT remains essentially on par with full fine-tuning (F-score $61.0$ vs.
+$61.6$), suggesting that merely keeping the “most sensitive” parameters does not yield a meaningful adaptation gain beyond standard tuning.
+This observation is consistent with the nature of multi-modal tracking: the adaptation signal is typically heterogeneous across layers and operators, and restricting updates to a subset can easily break the coordinated adjustments needed for modality fusion and temporal consistency.
+In contrast, our method achieves a substantially higher F-score ($65.1$) while keeping $\tau=100\%$ trainable parameters, demonstrating that the key factor is not sparsifying the update but steering it.
+By reweighting parameter changes according to their significance, our regularization encourages the model to preserve parameters that are critical to general tracking behavior while allowing more flexible updates on parameters that are more transferable for cross-modal adaptation.
+As a result, our method improves both precision and recall (Pr $64.7$, Re $65.4$), indicating more accurate target localization and more complete target recovery under depth-induced appearance variations.
+Table 13: Comparison of the tracking performance between SPT and our regularized fine-tuning method based on the DepthTrack dataset, using the pre-trained OSTrack-B256 as the base model.
+We have set up a series of trainable parameter ratios $\tau$ of SPT (top-$\tau$ sensitive parameters) to fully explore its adaption effect.
+Exp. Pr Re F-score
+w/o fine-tuning ($\tau=0\%$) 38.2 36.0 37.1
+full fine-tuning ($\tau=100\%$) 61.7 61.5 61.6
+SPT ($\tau=50\%$) 61.1 60.9 61.0
+SPT ($\tau=20\%$) 60.7 59.7 60.2
+SPT ($\tau=10\%$) 56.2 52.8 54.5
+Ours ($\tau=100\%$) 64.7 65.4 65.1
+Table 14: Computational cost of FIM eigen-decomposition on OSTrack-B256 ($\times128$$\times256$ input resolution for template and search regions).
+Metric Granularity Computational Cost
+Per-Group (Operation-wise) Per-Iteration (Model-wise) Total Offline Process
+FLOPs Theoretical Ops 104.8 GFLOPs 15.5 TFLOPs $\sim$ 39.7 EFLOPs
+Latency Wall-clock Time $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
+[…]
+Estimated over the full pre-trained tracking datasets. The preprocessing is performed only once.
+C.2 Computational Overhead of Prior Significance Estimation
+To theoretically justify the efficiency of prior significance estimation, we compare our approximation with the standard full-matrix estimation.
+Calculating the exact FIM on the global parameter space ($P\approx 86\text{M}$ for ViT-B) is computationally intractable, carrying a time complexity of $\mathcal{O}(P^{3})$ and a space requirement of $\mathcal{O}(P^{2})$ (theoretically requiring Petabytes of memory).
+In stark contrast, our SRFT adopts a group-wise approximation (operation-diagonal FIM) combined with Rayleigh quotient probing ($N_{\text{ops}}=148$ and $d=768$ for ViT-B).
+This strategy reduces the time complexity to $\mathcal{O}(N_{\text{ops}}\cdot d^{3})$—scaling linearly with the number of operations—and strictly bounds the peak memory to $\mathcal{O}(d^{2})$.
+Given that the operation width $d$ is magnitudes smaller than the total parameters $P$ ($d\ll P$), our method effectively reduces the computational cost from intractable to a practical level.
+We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table 14 to quantify its efficiency.
+While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior significance estimation introduces additional, non-trivial computation.
+Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference.
+Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle.
+C.3 Computational Overhead of Prior Significance under Larger Backbones
+In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior significance estimation under larger backbones, such as ViT-L and Swin-L.
+Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Table 15.
 It is important to note that the total cost of the offline FIM process is data-dependent, scaling with the size of the pre-training dataset and the number of iterations required for convergence given the model capacity.
 Consequently, rather than providing a variable total estimate, we report the deterministic per-group and per-iteration costs to accurately reflect the computational costs.
-TABLE XIV: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
+Table 15: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
 Model Metric Granularity Computational Cost
@@
 […]
-B-B Training Efficiency of SRFT
+C.4 Training Efficiency of Significance-Regularized Fine-Tuning
 We explicitly acknowledge that unlike plug-and-play PEFT methods (e.g., Prompt Tuning, Adapters, LoRA), which incur zero preprocessing cost via random initialization, our SRFT involves an offline FIM estimation phase.
@@
 Accelerated Convergence: The FIM estimation is a one-shot process performed on the pre-training datasets.
-The resulting sensitivity map is then frozen and reused for all subsequent multi-modal tracking tasks.
+The resulting significance map is then frozen and reused for all subsequent multi-modal tracking tasks.
 Furthermore, the pre-trained weight acts as a “smart initialization” compared to the random initialization of PEFT.
@@
 As shown in Tab.
-XV and Fig.
-11, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
+16 and Fig.
+12, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
 2.
@@
 Superior Performance: Finally, the computational cost translates directly into performance gains.
-By leveraging the global sensitivity information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
-TABLE XV: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods.
+By leveraging the global significance information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
+Table 16: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods.
 All experiments are conducted based on the pre-trained OSTrack-B256.
-“Preprocessing” refers to prior sensitivity estimation.
+“Preprocessing” refers to prior significance estimation.
 The training epochs (listed sequentially from left to right) correspond to the FE108, VisEvent, CoeSot, LasHeR, and DepthTrack datasets, respectively.
@@
 SDSTrack + SRFT 92.4 47.8 h $40$ $19.4min\times 40\approx 10.3\text{ h}$ 58.1
+C.5 Different Significance Fusion Scheduling Strategies
+Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain).
+In the paper, we explicitly define a dynamic linear schedule to harmonize the two significance: at the beginning, the prior significance has weight $\kappa$, and as training proceeds the transfer significance gradually increases until they reach the same weight $\kappa$ at the end.
+This is formalized by Eq.
+(14), where the combined significance is a convex combination with linear dependence on training progress $\frac{t}{T}$.
+As summarized in the Fig.
+13 and Tab.
+17, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under significance harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$.
+To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training.
+Overall, the linear schedule is consistently best or tied-best.
+When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime.
+However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile.
+Mechanistically, exponential/cosine may over-emphasize one significance at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization.
+In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
 Figure 12: Comparison of training loss for different multi-modal tracking methods on DepthTrack dataset.
@@
 Moreover, the training dynamics are noticeably more stable, exhibiting reduced oscillation and improved consistency throughout fine-tuning.
-Figure 12: Schematic illustration of different sensitivity fusion scheduling strategies.
-We compare the evolution of the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
-B-C Different Sensitivity Fusion Scheduling Strategies
-Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain).
-In the paper, we explicitly define a dynamic linear schedule to harmonize the two sensitivities: at the beginning, the prior sensitivity has weight $\kappa$, and as training proceeds the transfer sensitivity gradually increases until they reach the same weight $\kappa$ at the end.
-This is formalized by Eq.
-(14), where the combined sensitivity is a convex combination with linear dependence on training progress $\frac{t}{T}$.
-As summarized in the Fig.
-12 and Tab.
-XVI, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under sensitivity harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$.
-To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training.
-Overall, the linear schedule is consistently best or tied-best.
-When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime.
-However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile.
-Mechanistically, exponential/cosine may over-emphasize one sensitivity at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization.
-In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
-TABLE XVI: Comparison of different sensitivity fusion schedules on FE108 dataset.
-We compare the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
-Schedule Prior Sensitivity Weight Transfer Sensitivity Weight $\kappa$ SR PR
+Figure 13: Schematic illustration of different significance fusion scheduling strategies.
+We compare the evolution of the prior significance fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial significance harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
+Table 17: Comparison of different significance fusion schedules on FE108 dataset.
+We compare the prior significance fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial significance harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
+Schedule Prior Significance Weight Transfer Significance Weight $\kappa$ SR PR
 Exponential $w^{p}=\kappa\left(\frac{1-\kappa}{\kappa}\right)^{\frac{t}{T}}$ $w^{t}=1-w^{p}$ 0.3 65.7 94.2
@@
 0.9 66.9 95.5
-Figure 13: Visual comparisons of the tracking performance of different methods on the (Left) RGB-Event, (Middle) RGB-Depth and (Right) RGB-Thermal datasets.
-B-D Visual Comparison
+Appendix D Visual Comparison
 Some representative visualization results are illustrated in Fig.
-13.
+14.
 For clarity, all visual instances are explicitly labeled by the predicted bounding box.
@@
 These results demonstrate that our method improves tracking accuracy and provides a more stable, efficient way to integrate auxiliary data.
+Figure 14: Visual comparisons of the tracking performance of different methods on the (Left) RGB-Event, (Middle) RGB-Depth and (Right) RGB-Thermal datasets.